### v1

# Toward High-Performance LLM Serving: A Simulation-Based Approach for Identifying Optimal Parallelism

Abstract.
Large Language Models (LLMs) have achieved significant success in various domains, and serving LLMs efficiently has become crucial. LLMs are often served with multiple devices using parallelism techniques like data, pipeline, and tensor parallelisms. Each parallelism presents trade-offs between computation, memory, and communication overhead, making it challenging to determine the optimal parallel execution plan. Moreover, input workloads also impact parallelism strategies. Tasks with long prompts like article summarization are compute-intensive, while tasks with long generation lengths like code generation are often memory-intensive; these differing characteristics result in distinct optimal execution plans. Since searching for the optimal plan via actual deployment is prohibitively expensive, we propose APEX, an LLM serving system simulator that efficiently identifies an optimal parallel execution plan. APEX performs dynamism-aware simulation, which captures the complex characteristics of iteration-level batching, a technique widely used in state-of-the-art LLM serving systems. APEX leverages the repetitive structure of LLMs to reduce design space, maintaining a similar simulation overhead, even when scaling to trillion scale models. APEX supports a wide range of LLMs, device clusters, etc., and it can be easily extended to new models or clusters through its high-level templates. We run APEX simulations using a CPU and evaluate the optimal parallel execution plans found by APEX using a cluster with 8 H100 GPUs. The simulations cover a myriad of LLM architectures, precision formats, cluster configurations, and input workloads. We show that APEX can find optimal execution plans that are up to 4.42$\times$ faster than heuristic plans in terms of end-to-end serving latency. APEX also reports a set of metrics used in LLM serving systems, such as time per output token (TPOT) and time to first token (TTFT). Furthermore, APEX can identify an optimal parallel execution plan within 15 minutes using a CPU. This is 71$\times$ faster and 1234$\times$ more cost-effective than actual deployment on a GPU cluster using cloud services. APEX will be open-sourced upon acceptance. Finally, we show that LLM service providers can utilize APEX to meet service-level objectives and explore hardware design space to build high-performance LLM serving platforms.
LLM Serving, Optimal Parallelism, Simulation
copyright: acmlicensedjournalyear: 2018doi: xxx.xxxconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Modeling and simulationccs: Computing methodologies Parallel computing methodologies
1. Introduction
Large Language Models (LLMs) have been successfully applied to various applications, such as code generation (Fakhoury et al., 2024; Zhong and Wang, 2024), question-answering (Zhuang et al., 2023; Lucas et al., 2024), and many more (Zhong et al., 2023; Wang et al., 2024; Zhou et al., 2024). Serving an LLM is both memory- and compute-intensive; therefore, leveraging a cluster with multiple devices is essential to gain computing power and memory resources for achieving high performance. Several techniques have been proposed to parallelize LLM on multiple devices, such as data parallelism (DP) (Li et al., 2020b), pipeline parallelism (PP) (Huang et al., 2019), tensor parallelism (TP) (Shoeybi et al., 2020), etc. Each parallelism has its pros and cons. Techniques like TP are more memory-efficient than DP as it does not produce model replicas, but they incur expensive collective communication overhead (Shoeybi et al., 2020). On the other hand, PP has lower communication overhead but can suffer from workload imbalance (Huang et al., 2019). To balance the trade-offs, parallelism techniques can be adopted in a hybrid manner, potentially leading to better serving performance than relying on a single type of parallelism. However, determining the optimal parallel execution plan is challenging due to the many factors involved, including computation workload, network traffic, memory efficiency, etc. Furthermore, the optimal parallel execution plan also depends on the characteristics of the input requests. Some requests lead to long prompts with short generation lengths, such as summarizing an article, while some lead to short prompts and long generation lengths, like code generation and storytelling. The former type of request is compute-intensive, while the latter is memory-intensive. As a result, these two types of requests favor different parallelism techniques to optimize performance. A common practice is to adopt heuristic execution plans instead, such as applying TP within the same node and PP across different nodes (Miao et al., 2024; vLLM, v060). Yet, a recent study has shown that applying such heuristics approaches can be up to 2$\times$ slower than the optimal configuration (Isaev et al., 2023); this calls for the need for service providers to search for and adopt an optimal parallel execution plan rather than relying on heuristics.
To find an optimal execution plan, a straightforward approach is to deploy and evaluate various parallel execution plans directly; however, this approach is prohibitively expensive, as it could take thousands of GPU hours to assess multiple execution plans (Agrawal et al., 2024b, a). Note that this high searching cost cannot be amortized, as the optimal configuration varies depending on the model and the workload characteristics of the input requests. An alternative is to search via modeling-based approaches, such as developing performance models (Isaev et al., 2023) or building simulators (Yu et al., 2021; Duan et al., 2024) to estimate the performance of a given parallel execution plan; nevertheless, existing solutions developed for Deep Neural Network (Yu et al., 2021; Duan et al., 2024) or LLM training (Isaev et al., 2023; Bang et al., 2024) cannot be applied to LLM serving, as such systems introduce unique challenges: (1) Dynamism of iteration-level batching: Unlike conventional ML systems that adopt static batching, which waits for all requests in the current batch to be completed before batching new requests, state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) adopt iteration-level batching (Yu et al., 2022) to achieve high serving performance. With iteration-level batching, incoming requests are continuously added to the processing batch whenever memory becomes available. This makes it challenging to model the system, as the batch size changes dynamically during serving. Additionally, some requests within the batch may be in the prefill stage, while others are in the generation stage. This interleaving of stages further complicates the modeling, as these stages have significantly different computational characteristics (details in Section 2.1). (2) Exponentially-growing design space: Although modeling-based approaches provide a potentially time-efficient solution to search for the optimal parallel execution plan, they can still incur substantial overhead as the design space grows exponentially with respect to the model size and the number of devices in the cluster. Since LLM serving (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) often involves large models and clusters, it is non-trivial to develop a solution that can search for an optimal parallel execution plan within a reasonable timeframe. (3) Adapting for continuously-evolving systems: LLM serving system is an emerging area, with continuous evolution in model architecture, hardware platforms, and system optimizations such as quantization (Lin et al., 2024; Hooper et al., 2024) and parallelism techniques (Rajbhandari et al., 2022; Liu et al., 2023). Consequently, a performance model or a simulator can easily become obsolete and inapplicable to state-of-the-art LLM serving systems. Thus, keeping pace with the rapid evolution of such systems is necessary but also challenging.
To this end, we propose APEX, an extensible and dynamism-aware simulator for automated parallel execution in LLM serving. APEX takes an LLM, a set of input requests with various context and generation lengths, and a device cluster as inputs, and generates an optimal parallel execution plan for LLM serving. Specifically, for a given LLM model and device cluster, APEX first generates various parallel execution plans, each representing a unique way to parallelize the model by combining various parallelism techniques, such as tensor, data, pipeline, and expert parallelisms. APEX then evaluates each plan by estimating the execution time of serving the input requests through simulation. APEX performs dynamism-aware simulation that is capable of modeling the complex characteristics of iteration-level batching, which involves concurrently serving requests of varying lengths and stages (i.e., prefill and generation stage). After the simulation, APEX provides a comprehensive evaluation for each parallel execution plan, which includes multiple metrics used in LLM serving systems, such as time per output token (TPOT), time to first token (TTFT), end-to-end serving latency, P95 latency, among others (Miao et al., 2023). A parallel execution plan with the lowest end-to-end serving latency is chosen as the optimal plan. Despite performing complex dynamism-aware simulations, APEX remains time-efficient by leveraging the repetitive structure of transformer layers in LLMs, which significantly reduces the design space; this allows APEX to scale to trillion-scale models on multi-node device clusters. APEX supports a broad range of LLMs, device clusters, quantization formats, and parallelism techniques, which are essential for modeling state-of-the-art LLM serving systems. Recognizing the rapid evolution of LLM serving systems, APEX is designed to be modular and extensible. We capture the high-level abstractions of LLMs and device clusters, and develop software templates based on these abstractions. This approach allows new models and device clusters to be easily supported with minimal coding effort in the templates. Additionally, APEX leverages operation-level profiling results to estimate the execution time of parallel execution plans on a given device cluster. Profiling-based estimation enables APEX to adapt to different clusters by collecting profiling data from the underlying platform. Note that the device cluster APEX supports is not limited to GPU clusters, but also AI-accelerator clusters such as TPU (Jouppi et al., 2023), Intel Gaudi (Intel, 2024), etc. We plan to open-source APEX upon paper acceptance. We further demonstrate that, in addition to identifying optimal parallel execution plans, APEX can also assist LLM service providers in meeting service-level objectives (SLOs) and provide valuable insights for building LLM-serving hardware platforms. The key contributions of this work are:
  • We propose APEX, a dynamism-aware simulator that captures the complex characteristics of LLM serving. APEX automatically finds an optimal parallel execution plan for a given LLM, a device cluster, and a set of input requests.
  • APEX supports a wide range of LLMs, parallelism techniques, quantization formats, and device clusters. APEX can also be easily extended to new models and device clusters.
  • We evaluate APEX using four LLMs on three datasets with distinct workloads. APEX identifies optimal parallel execution plans that are up to 4.42$\times$ faster than heuristic plans in terms of the end-to-end serving latency.
  • APEX provides high-fidelity simulation, achieving less than 10% relative error on average when predicting the speedup of optimal execution plans over the baseline plans.
  • APEX is time-efficient and cost-effective, capable of finding an optimal parallel execution plan within 15 minutes on a CPU; this is 71$\times$ faster and 1234$\times$ cost-effective than actual deployment on a GPU cluster using cloud services.
  • APEX is highly scalable, maintaining similar simulation overhead when scaling from billion-scale to trillion-scale models.
  • APEX features comprehensive evaluation, which can help service providers meet SLOs. APEX also provides insightful suggestions for building LLM-serving hardware platforms.
2. Background
2.1. LLM Inference
Large Language Models (LLMs) are a type of ML models that are built upon transformer architecture (Vaswani et al., 2017). LLM inference takes a user prompt as input, and generates a response, token by token, sequentially. The sequential token generation process is also known as the auto-regressive generation, which utilizes the previously generated tokens as input to predict the next token. The inference process of LLMs consists of two stages:
Prefill Stage: In the prefill stage, LLM processes the input request (i.e., prompt) to set up intermediate states (keys and values) that are used to predict the first token. Unlike token generation, the computation of prefill does not rely on previously generated output tokens, allowing the tokens of the input request to be processed in parallel all at once. This high degree of parallelism makes the prefill stage compute-bound (Zhong et al., 2024).
Generation Stage: During the generation phase, LLM generates output tokens autoregressively until a stopping criterion is met. The generation of each token depends on all previous tokens’ output states (keys and values). The generation phase is often memory-bound (Lin et al., 2024), as the latency mainly depends on the speed of data transfer for the output states from the memory, rather than on the computation itself.
2.2. LLM Serving and Iteration-Level Batching
LLMs are often hosted by service providers in the cloud (Li et al., 2024). Users send requests to these providers through APIs or chatbots and then receive responses generated by the LLMs. While a single LLM inference processes a single input request and generates one output response, LLM serving concurrently processes multiple input requests from the users and generates multiple output responses, i.e., multiple LLM inferences happen in parallel. For LLM serving, it is critical to achieve both low latency and high throughput: low latency is necessary to meet service-level objectives, while high throughput enables service providers to serve a large number of users simultaneously. Due to the autoregressive nature of LLMs, the lengths of the generated responses can vary significantly. Consequently, the commonly used static batching leads to suboptimal performance in LLM serving, as all the batched requests need to wait for the longest response to be generated to proceed to process new requests. To overcome this inefficiency, iteration-level batching (Yu et al., 2022) is proposed. Iteration-level batching continuously schedules newly arrived request(s) into the existing batch whenever GPU memory becomes available, rather than waiting for all the requests to be completed. Iteration-level batching significantly improves the serving throughput and is widely adopted in state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024).
2.3. Device Clusters and Interconnections
Serving an LLM is both memory- and compute-intensive. To achieve high serving throughput and accommodate the large model size, LLMs are often deployed on a cluster of devices. The devices are connected in a tree-based network topology, which is one of the most popular topologies for node deployment. The memory bandwidth and latency is uniform within each level. Figure 1 illustrates an example of a commonly used two-level device cluster. Devices within the same node are typically connected by high-bandwidth interconnects like NVLink (Li et al., 2020a) at level 1, while devices across different nodes are connected via inter-node networks such as InfiniBand (Liu and Guitart, 2022) at level 2. In addition to GPUs, clusters of AI accelerators are also emerging, such as TPU clusters (Jouppi et al., 2023) and Gaudi clusters (Intel, 2024). These clusters also utilize tree-based network topologies for interconnection and can be abstracted similarly as GPU clusters (Choquette, 2023).
Figure 1. An example of a two-level device cluster hierarchy, where leaves represent the devices. The memory bandwidth and latency are uniform within the same level.
Figure 2. System overview of APEX
2.4. LLM Parallelisms
LLM is often served with multiple devices (Figure 1), and various approaches have been proposed to parallelize LLM on the cluster. We discuss several representative and widely-used parallelisms.
Data Parallelism (DP): In DP (Li et al., 2020b), the model is replicated across multiple devices. The input requests are split into micro-batches and distributed to each model replica for processing. DP incurs no communication overhead as each micro-batch is processed independently. However, having model replicas incurs large memory overhead, and reduces the number of requests that can be batched concurrently.
Pipeline Parallelism (PP): PP (Huang et al., 2019; Narayanan et al., 2019) divides the model in a layer-wise fashion. Each model partition consists of a subset of layers, and each subset of layers forms a pipeline stage. The input requests are split into micro-batches to flow through the pipeline stages. PP requires point-to-point (p2p) communications between the pipeline stages. In addition, PP suffers from pipeline stalls when the execution time of each stage is unbalanced.
Tensor Parallelism (TP): TP (Shoeybi et al., 2020) divides the model in an intra-layer fashion, which splits individual layers across multiple devices. TP does not suffer from load imbalance like PP or produces model replicas like DP. Yet, TP incurs high collective communication overhead, such as AllReduce, which is prohibitively expensive for interconnection networks with limited bandwidth like PCIe.
Expert Parallelism (EP): EP (Rajbhandari et al., 2022) is a special type of parallelism for Mixture of Experts (MoE) models. EP deploys different experts across various devices. For MoE models, only a subset of experts are activated for each input token. Thus, EP can result in workload imbalance and resource under-utilization if none of the experts on the device are activated; while this can be avoided by adopting TP, where each device is assigned with a partition of all the experts, EP incurs a lower communication overhead than TP.
Each parallelism has its own pros and cons, tradeing-off between computation, memory efficiency, collective communication overhead, etc. Thus, it’s non-trivial to determine the optimal parallel execution plan, especially given that the parallelisms can be adopted in a hybrid manner. Heuristic plans are typically adopted for simplicity, such as applying TP to clusters with high-bandwidth interconnections like NVLink (Li et al., 2020a), and applying PP to clusters without such networks (Miao et al., 2024; vLLM, v060). APEX aims to evaluate the complex trade-offs among parallelism techniques to find optimal parallel execution plans that outperform heuristic approaches.
2.5. LLM Quantizations
Quantization is an essential technique for achieving performant LLM serving, as it reduces both memory and computation overhead. Various methods have been proposed to quantize one or more of the following components to lower precision while preserving high accuracy: (1) model weights, (2) activations, and (3) KV cache. For example, AWQ (Lin et al., 2024) quantizes the model weights, and KVQuant (Lin et al., 2024) focuses on quantizing the KV cache. State-of-the-art LLM serving systems like vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2024) also support W8A8 quantization, which quantizes both the model weights and activations to FP8 format. Given the diversity of quantization methods, it is essential to flexibly support various techniques to find an optimal parallel execution plan for LLM serving systems.
3. APEX Simulator Design
3.1. System Overview
We depict the system overview of APEX in Figure 2. Initially, the Offline Profiler (Section 3.2) of APEX takes a set of profiling scripts to obtain the performance information of the underlying platform, and the results are stored for APEX’s profiling-based simulation. Profiling is an offline process that is only performed once when porting to a new cluster with unknown devices (e.g., porting from an A100 to an H100 GPU cluster.) Next, given an LLM and device cluster, the Parallel Execution Plan Generator converts the model into an intermediate representation (Section 3.3) and generates various parallel execution plans (Section 3.4 and 3.5); each plan represents a unique way to map the LLM onto the device cluster. The Iteration-Level Batching Module (Section 3.6) takes the generated parallel execution plans and a set of request traces as inputs, and starts batching the input requests. The Batching Module determines whether a request should be added to or removed from the batch in each iteration, and reports the active requests to the Dynamism-Aware Simulator (Section 3.7). Based on the active requests in the batch, the Dynamism-Aware Simulator estimates the execution time for each iteration and keeps track of multiple LLM system metrics such as time to first token (TTFT), time per output token (TPOT), P95 latency, etc. After all requests have been processed, the Dynamism-Aware Simulator produces a Simulation Report for the corresponding parallel execution plan. This process is repeated for all the generated parallel execution plans, and the plan with the lowest end-to-end latency is selected as the optimal plan.
3.2. Offline Profiler
The Offline Profiler collects performance information of the underlying platform, where the results are used by the Dynamism-Aware Simulator (Section 3.7) to estimate the execution time of different parallel execution plans. The Offline Profiler leverages the fact that LLMs are based on transformer architecture, which consists of similar operations, and performs operation-level profiling, such as measuring the computation time of multi-head attention, which is used in the attention layers, and general matrix multiplication (GEMM), which is used in the feedforward layers. Profiling the key transformer operations allows APEX to support various LLMs without exhaustively profiling each model, as they can be broken down into the same set of operations. In addition to computation operation, the Offline Profiler also profiles the collective communication overheads on the cluster, e.g., the time to perform AllReduce, ReduceScatter, All-to-All. This allows APEX to support various parallelisms, which require different collective communications. For the computation operations, the profiling is performed by going through various sequence lengths, number of attention heads, hidden dimensions, etc. For the collective communication operations, the profiling is performed by varying the data transfer sizes and also the number of devices involved, both within a node and across nodes, in the collective operation. Operation-level profiling produces profiling results that can support various LLMs, request traces, parallelisms, and device clusters of different sizes. The profiling only needs to be performed once when porting to a new cluster with unknown devices, such as porting from an A100 GPU cluster to a H100 GPU cluster. Therefore, for the same type of device cluster, the profiling overhead is a one-time cost that can be amortized.
Figure 3. Transformer IR represents LLMs in a canonical way
3.3. Transformer IR
APEX identifies an optimal parallel execution plan by generating and evaluating multiple candidate execution plans. Yet, it is non-trivial to generate parallel execution plans for a variety of LLMs as they have different model architectures. To address this, we introduce the Transformer IR, a unified abstraction for transformer-based models. APEX utilizes Parallel Templates, developed based on the Transformer IR rather than specific models, to generate parallel execution plans. This approach enables APEX to parallelize a wide range of LLMs that can be represented through the Transformer IR. We discuss the details of Parallel Templates in Section 3.4.
We depict the idea of Transformer IR in Figure 3. We define the key operations of transformers, such as the multi-head attention, as cells. An LLM can be represented as a chain of cells. For example, GPT-3 (Brown et al., 2020) models can be represented as a chain of multi-head attention cells and multi-layer perceptron cells, and Llama-3.1 (Dubey et al., 2024) models can be represented as a chain of group query attention (Ainslie et al., 2023) cells and SwiGLU (Shazeer, 2020) cells. The Transformer IR represents an LLM canonically, which only captures the key cells in the model and ignores operations like tokenization and position embeddings, as they are less relevant for model parallelization; this reduces the search space for simulation. Given that LLMs often use the same set of cells repeatedly, we further define the smallest set of non-repetitive adjacent cells as a block. Each cell consists of multiple tasks, such as the attention heads in the multi-head attention cell, or the experts in the mixture-of-expert cell. Each task works independently from the other tasks (i.e., no inter-communication is needed), and their outputs are combined via operations like concatenation or AllReduce. To summarize, using the Transformer IR, an LLM can be unifiedly represented as multiple identical blocks, each block consists of numerous cells, and each cell consists of multiple tasks. We develop an IR converter, which parses the information of an LLM configuration file, such as the number of attention heads, number of layers, etc., and represents the model using Transformer IR.
3.4. Parallel Templates
We develop Parallel Templates to generate various parallel execution plans for LLM serving. As mentioned in Section 3.3, the templates are developed based on the Transformer IR, which allows the templates to support a wide range of LLMs. Each cell type is associated with a pre-defined Parallel Template that describes how it can be parallelized on multiple devices. We show some examples in Figure 4. If tensor parallelism is applied to both the multi-head attention (MHA) cell and the multi-layer perceptron (MLP) cell, as in Figure 4 (a), the tasks (e.g., attention heads) in the two cells are evenly distributed to each device. In addition to distributing the tasks, the Parallel Templates also handle the collective communications between adjacent cells. In the example of Figure 4 (a), an AllReduce is performed between the cells for data synchronization, which is similar to the case in Megatron-LM (Shoeybi et al., 2020). If the numbers of cell replicas vary between the adjacent cells, tensor resharding is required for the output activation; such cases happen when different data parallelism (DP) degrees are applied. For example, in Figure 4 (b), a 2-way DP is applied to the MLP cell, and a 1-way DP (i.e., no DP) is applied to the MHA cell, resulting in a different number of cells replicas. Consequently, All-to-All and AllGather operations are performed for tensor resharding. Figure 4 (c) depicts a Parallel Template associated with the MHA cell: $H$ attention heads are evenly distributed to $D$ device for computation, and collective communications for resharding and synchronization may also be performed, depending on the number of cell replicas of the adjacent cells. Note that while Figure 4 shows a simplified example of having only two devices, the Parallel Templates are parameterized and can support any number of devices.
Figure 4. Examples of APEX’s Parallel Templates for two devices. The parameterized templates can extend to D devices.
3.5. Parallel Scheme Generator & Device Mapper
Given an LLM and pre-defined Parallel Templates, the Parallel Scheme Generator produces various parallel schemes, which maps the LLM onto a logical device cluster (Zhao et al., 2020) with various combinations of parallelism techniques. A logical device cluster is a virtual overlay that describes the number of devices in the cluster without the information of the network topology. Consequently, a parallel scheme only defines the parallelisms applied to the Transformer IR cells and their corresponding collective communications, such as the examples in Figure 4 (a) and (b), without assigning the cells to specific physical devices. We introduce a virtual overlay of a logical device cluster for design simplicity; this approach allows us to decouple the Parallel Template designs (Section 3.4) from the hierarchal network topology of a device cluster - jointly considering the network topology in the Parallel Templates would greatly increase design complexity. The Parallel Scheme Generator creates parallel schemes using a hierarchical top-down approach, beginning with the most coarse-grained model-level parallelism and gradually refining to fine-grained cell-level parallelism. The detailed procedure is outlined in Algorithm 1. First, the Generator decides the degree of model-level data parallelism (i.e., the number of model replicas), and also derives the number of devices assigned to each replica (i.e., number of replica devices). APEX requires the device cluster to be evenly partitioned by the number of model replicas, so the available degrees of parallelism are restricted to the divisors of the total number of devices in the cluster. Second, the Generator decides the number of pipeline stages within each replica. Similarly, APEX requires the cluster to be evenly partitioned by the number of stages, so the available parallelism degrees are the divisors of the number of devices assigned to the replica. Third, the Generator decides the degree of cell-level data parallelism for each cell, and generates task mappings using the pre-defined Parallel Templates. If the degree of cell-level data parallelism for a specific cell is less than the number of assigned devices, the Parallel Templates apply intra-layer parallelism like tensor or expert parallelism to parallelize the cell (Section 3.4). For example, if four devices are assigned to a cell with two replicas, two devices are assigned to each replica, and the cell can be parallelized using tensor parallelism. Afterward, for each pair of adjacent cells, the Generator inserts collective communications into the parallel schemes if tensor resharding is needed.
Algorithm 1 Parallel Scheme Generation
1:Input: LLM (in Transformer IR), Parallel Templates
2:Output: parallel_schemes
3:$n$ = num_of_devices_in_cluster
4:for $model\_DP=1,2,...,n$do $\triangleright$ model-level DP
5:     $m$ = $n\div model\_DP$triangleright $m$ = num_repica_devices
6:     for $stage=1,2,..,m$ do $\triangleright$ inter-layer parallelism
7:         $s$ = $m\div stage$triangleright $s$ = num_stage_devices
8:         for cell $\in$ LLM.block do
9:              for $cell\_DP=1,2,...,s$do $\triangleright$ cell-level DP
10:                  $c$ = $s\div cell\_DP$triangleright $c$ = num_cell_devices
11:                  task_mapping = templates(cell, c)
12:                  cell_schemes.append(task_mapping)
13:         for $r$ in 0,…, len(LLM.block.cells): do
14:              reshard = get_reshard_collective(cell_schemes, $r$)
15:              stage_schemes.append(cell_schemes, reshard)
16:         parallel_schemes.append(stage_schemes)
After generating the parallel schemes, the Device Mapper maps the logical devices onto the physical device cluster, producing parallel execution plans. While the parallel schemes are generated in a top-down approach, the Device Mapper operates in a bottom-up manner. The logical devices are first mapped to physical devices connected at the bottom level of the cluster (see Figure 1); if the number of logical devices exceeds the available physical devices at the current level, the Device Mapper moves to the next upper level to include additional physical devices for mapping. Since lower-level device connections generally offer higher bandwidth than upper-level connections, the Device Mapper prioritizes mapping logical devices assigned to the same cell, as fine-grained cell-level parallelisms tend to incur expensive collective communication overhead, such as the AllReduce in tensor parallelism. The Device Mapper then maps logical devices assigned to the same pipeline stage, followed by those assigned to the same model replica, progressing through increasingly coarse-grained parallelism levels. This bottom-up approach maps logical devices with potentially higher communication overhead to physical devices connected at the lower level, which have higher interconnection bandwidth, and vice versa.
3.6. Iteration-Level Batching Module
The Iteration-Level Batching Module takes the candidate execution plans and request traces as inputs, and simulates the behavior of a serving system using iteration-level batching (Yu et al., 2022). Each request comprises three key attributes: context length, generation length, and arrival time. As mentioned in Section 2.2, iteration-level batching continuously adds arriving requests to the current batch whenever memory permits. Therefore, a Resource Monitor keeps track of the memory usage and determines whether a new request can be batched. The Request Scheduler also checks the arrival time of a request to determine whether a request has arrived and can be batched. The Request Scheduler maintains a list of active requests, which are the requests in the current batch. The active request list also tracks how many tokens have been generated for each request. If the number of generated tokens matches the generation length of a request, then the request is completed and should be removed from the active batch; this releases memory to batch new requests. One token is generated for each active request in one iteration. In each iteration, the Iteration-Level Batching Module sends the active request list to the Dynamism-Aware Simulator (Section 3.7) to estimate the execution time. The Request Scheduler greedily adds arriving requests to the batch whenever memory is available to accommodate the context length of a request, without considering the memory required for the KV cache of generated tokens. As a result, memory capacity may be exhausted before token generation for all requests is completed. In such cases, the most recently added requests and their generated tokens are temporarily removed from the batch to free up memory for earlier requests to complete their token generation. These removed requests are re-added to the batch as memory becomes available. Both the removal and re-addition are performed in order, prioritizing completing the earliest requests.
3.7. Dynamism-Aware Simulator
Given the active requests produced by the Iteration-Level Batching Module and the Profiling Results, the Dynamism-Aware Simulator estimates the execution time of each iteration. The active request list contains information such as the context length of the request and the number of tokens that have already been generated for a request. This allows the Simulator to identify whether a request is in the prefill stage or the generation stage. Specifically, a request is in the prefill stage if the number of generated tokens equals zero; otherwise, it is in the generation stage. For requests in the prefill stage, we calculate the execution time of the operations by setting the input sequence length as the context length, as the entire sequence is processed in parallel (Section 2.1). For requests in the generation stage, we calculate the execution time of the operations by setting the input sequence length as one; this is because the output states (i.e., key and value) of the previous tokens are stored in the KV cache and do not require recomputation. Thus, the computation is only performed on the current token, reducing the effective sequence length to one. Furthermore, requests in the generation stage are processed in parallel; thus, assuming $n$ generation requests, the Simulator estimates the execution time by setting the sequence length as $n$, as each request has an effective sequence length of one. The Profiling Results contain the execution time of different attention heads, hidden dimensions, sequence lengths, etc., allowing the Simulator to estimate the total execution time for processing the prefill requests and the generation requests. If a specific data point is not presented in the Profiling Results, the Simulator estimates its value by leveraging linear interpolation between the nearest available profiling points. The Simulator is dynamism-aware as the Iteration-Level Batching Module continuously updates the list of active requests in every iteration. In addition, the Simulator utilizes the repetitive nature of LLMs to reduce simulation overhead. As discussed in Section 3.3, a block in Transformer IR is the smallest set of non-repetitive cells. Thus, the Simulator performs simulation using a single block and projects the execution time of the entire model, which consists of multiple repetitive blocks. The projection calculation, such as summing or taking the maximum of the execution times, depends on the connection pattern of the blocks (e.g., sequential or pipelined). In addition to estimating the execution time of each iteration, the Simulator also features a Metric Monitor, which calculates the value of several important metrics in LLM serving systems, such as time to first token (TTFT), time per output token (TPOT), P95 latency, Model FLOPs Utilization (MFU), Model Bandwidth Utilization (MBU), among others.
4. Experimental Results
Table 1. Detail of the request traces used for evaluation
Traces Context Lengths Generation Lengths # of requests
Summarization 2742.11$\pm$944.33 172.22$\pm$73.17 1188
Creation 306.82$\pm$81.03 1128.34$\pm$419.64 512
Chat 73.32$\pm$148.65 189.47$\pm$174.18 1024
4.1. Experimental Setup
To evaluate APEX, we choose a broad range of LLMs and datasets with distinct workloads. We also experiment with different data types. We discuss the setup details below.
Models: For evaluation, we choose four state-of-the-art LLMs: Qwen2.5-32B-Instruct (Hui et al., 2024), Llama-3.1-70B-Instruct (Dubey et al., 2024), Mistral-Large-Instruct (123B) (Jiang et al., 2023), and Mixtral 8x22B (Jiang et al., 2024); This covers a myriad of LLMs in different sizes, ranging from 32B to 123B, and also a Mixture-of-Expert (MoE) model. While the models use half-precision (FP16) by default, we also test cases where the model is quantized to FP8 format.
Datasets: We prepared three sets of request traces derived from distinct datasets: a paper abstract summarization dataset (Cohan et al., 2018), a news abstract summarization dataset (Fabbri et al., 2019), and a conversational dataset, LMSYS-Chat-1M (Zheng et al., 2024). These datasets represent three distinct workloads. The paper abstract summarization (Cohan et al., 2018) exemplifies a prefill-intensive workload, which has a long context length and short generation length. While some LLM workloads, such as paper summarization, are prefill-intensive, some (e.g., code generation and storytelling) are generation-intensive, involving short context lengths and long generation lengths. To create a generation-intensive workload, we adapt the news abstract summarization dataset (Fabbri et al., 2019), which consists of short summarizations. We modified the dataset by appending the following prompt to each news summarization: “Please generate a long story using the provided abstract,” and use them as input. This prompts the LLM to produce a long story from a short abstract, resulting in a generation-intensive workload. The LMSYS-Chat-1M (Zheng et al., 2024) dataset collects a series of real-world conversations between users and LLMs. Most requests in this dataset feature short context and generation lengths, representing a lightweight conversational workload. For evaluation, we randomly subsampled between 512 and 1K requests from each dataset. To simplify terminology, we refer to these datasets as Summarization (paper abstracts), Creation (news generation), and Chat (LMSYS-Chat-1M) in the following sections. We assume a Poisson distribution for request arrival times. Details of the request traces are listed in Table 1. We report context lengths and generation lengths in terms of tokens, which vary for different tokenizers. We use the Llama-3.1 tokenizer as an example in Table 1; other tokenizers produce similar token counts.
Hardware Platform and Serving Systems: We run APEX simulation using an Intel Xeon 6530 CPU and validated the simulation results against the actual performance of LLM serving on a GPU cluster. We deployed vLLM (Kwon et al., 2023) v0.6.0 to configure an LLM online server on a GPU cluster equipped with 8 Nvidia H100 SXM GPUs, each with 80 GB of GPU memory. Requests were sent to the server according to their arrival times. Since vLLM does not natively support data parallelism, we implemented data parallelism by setting up multiple vLLM servers and dispatching requests to them in a round-robin fashion.
Table 2. APEX simulation results: End-to-end latency (seconds). APEX demonstrates its ability to identify optimal parallel execution plans, achieving superior serving performance compared to heuristic baseline plans.
Traces Model Arrival Rate Baseline Feasible Optimal APEX Optimal
Summarization Qwen-2.5-32B 0.25 2153.27 (1$\times$) 1137.70 (1.89$\times$) 968.26 (2.22$\times$)
0.5 1823.85 (1$\times$) 1116.84 (1.63$\times$) 967.23 (1.89$\times$)
Llama-3.1-70B 0.25 1998.82 (1$\times$) 1340.99 (1.49$\times$) 1175.09 (1.70$\times$)
0.5 1565.11 (1$\times$) 1225.38 (1.28$\times$) 1140.19 (1.37$\times$)
Mistral-Large 0.25 1301.27 (1$\times$) 498.32 (2.61$\times$) 376.13 (3.46$\times$)
0.5 1140.36 (1$\times$) 498.32 (2.29$\times$) 376.13 (3.03$\times$)
Mixtral-8x22B 0.25 1841.90 (1$\times$) 1290.97 (1.43$\times$) 551.94 (3.34$\times$)
0.5 1429.46 (1$\times$) 1137.87 (1.26$\times$) 552.28 (2.59$\times$)
Creation Qwen-2.5-32B 0.25 5718.22 (1$\times$) 2432.1 (2.35$\times$) 1507.68 (3.79$\times$)
0.5 4950.96 (1$\times$) 2421.82 (2.04$\times$) 1512.95 (3.27$\times$)
Llama-3.1-70B 0.25 6549.85 (1$\times$) 3742.77 (1.75$\times$) 3159.96 (2.07$\times$)
0.5 5240.78 (1$\times$) 4578.49 (1.14$\times$) 3183.67 (1.65$\times$)
Mistral-Large 0.25 6379.09 (1$\times$) 4063.11 (1.57$\times$) 2684.37 (2.38$\times$)
0.5 5221.52 (1$\times$) 4121.53 (1.27$\times$) 2682.05 (1.95$\times$)
Mixtral-8x22B 0.25 2645.24 (1$\times$) 2423.82 (1.09$\times$) 1039.87 (2.54$\times$)
0.5 1442.15 (1$\times$) 1417.75 (1.02$\times$) 1035.62 (1.39$\times$)
Chat Qwen-2.5-32B 0.25 1251.75 (1$\times$) 798.75 (1.57$\times$) 513.38 (2.44$\times$)
0.5 1115.01 (1$\times$) 786.08 (1.42$\times$) 510.72 (2.18$\times$)
Llama-3.1-70B 0.25 1622.48 (1$\times$) 1118.85 (1.45$\times$) 824.43 (1.97$\times$)
0.5 1344.42 (1$\times$) 1021.09 (1.32$\times$) 808.96 (1.66$\times$)
Mistral-Large 0.25 973.16 (1$\times$) 396.32 (2.46$\times$) 254.26 (3.83$\times$)
0.5 890.71 (1$\times$) 391.96 (2.27$\times$) 254.26 (3.50$\times$)
Mixtral-8x22B 0.25 1757.68 (1$\times$) 1504.88 (1.17$\times$) 397.93 (4.42$\times$)
0.5 1406.75 (1$\times$) 1253.22 (1.12$\times$) 396.80 (3.55$\times$)
4.2. Evaluation of APEX’s suggestion
In this experiment, we address the following research question: Can APEX improve LLM serving performance by identifying an optimal parallel execution plan? For evaluation, we design multiple tasks, with each task comprising an LLM, request traces, and an arrival rate. We run APEX simulation for each task and compare the performance of the following three execution plans:
  • Baseline plan: We follow the commonly-used heuristics (Miao et al., 2024; vLLM, v060), which applies tensor parallelism within the same node and pipeline parallelism across nodes. Since we only run on a one-node cluster, we apply tensor parallelism as the baseline plan.
  • APEX Optimal plan: APEX identifies an optimal plan for each task. However, APEX’s search space extends beyond the capabilities of current LLM serving systems, including advanced features like cell-level data parallelism. As a result, an execution plan may not be fully supported by existing LLM serving systems.
  • Feasible Optimal plan: This is the optimal plan identified by APEX under the constraint that only parallelism techniques supported by current LLM serving systems are used. Due to the reduced search space, the feasible optimal plan may achieve lower performance compared to the unconstrained APEX Optimal plan.
While the APEX Optimal plan may not be realizable in current LLM serving systems, we include the results to showcase the potential performance gains that can be achieved by supporting a broader range of parallelism techniques.
We show the experimental results in Table 2. We experiment with the three request traces and four LLMs mentioned in Section 4.1. We also conduct experiments using two different arrival rates for the requests, assuming a Poisson distribution. To explore different quantization formats, we quantize the Mistral-Large model using FP8 for the KV cache, as well as the weights and activations (i.e., W8A8). APEX identifies optimal parallel execution plans for both dense and sparse LLMs (i.e, MoE models), and the Feasible Optimal plans consistently deliver performance improvement over the baseline, achieving up to 2.61$\times$ speedup in terms of end-to-end serving latency. Furthermore, assuming cell-level data parallelism is available, the APEX Optimal Plans can achieve a speedup of up to 4.42$\times$. Cell-level data parallelism is particularly effective when the execution time varies among cells (e.g., MHA and MLP cells). This approach allows a specific cell to be parallelized further using data parallelism without requiring replication of the entire model. In this experiment, we demonstrate APEX is able to identify an optimal parallel execution plan that improves LLM serving performance under various LLMs and request traces.
Figure 5. We compare the speedup predicted by APEX against the actual speedup achieved using vLLM. APEX accurately predicts the speedup of adopting an optimal execution plan over the baseline plan.
Figure 6. APEX accurately predicts the scalability trend across varying numbers of GPUs, as measured by TPOT (ms)
APEX identifies an optimal parallel execution plan for each task. The optimal plans consist of various combinations of data, pipeline, and tensor parallelism, effectively balancing the trade-offs of various factors (e.g., compute, memory, network traffic) to outperform the baseline plan, which only relies on tensor parallelism. While the optimal plans are different from task to task, we observe that incorporating data parallelism (DP) often yields performance benefits. Many of the identified optimal plans set the degree of DP to 2 or even 4. Existing LLM serving systems often overlook DP, assuming it is prohibitively memory-intensive. Instead, our results demonstrate that trading memory efficiency for reduced communication overhead can lead to performance improvements. We also observe that the Feasible Optimal plan of Mixtral-8x22B shows marginal improvement over the baseline on the Creation dataset. This is due to the dominating role of memory utilization in this setup, leaving fewer trade-offs available for performance improvement. Specifically, Mixtral-8x22B, with its 141B parameters, heavily utilizes memory resources, while the Creation dataset’s generation-intensive nature further increases memory demands due to extensive KV caching. Under these conditions, the baseline plan, which adopts pure tensor parallelism, performs well owing to its high memory efficiency. Nevertheless, the APEX Optimal plan still manages to deliver a 2.54$\times$ speedup by enabling cell-level data parallelism (DP), which is more memory-efficient than replicating the entire model.
4.3. Evaluation of Simulation Fidelity
In this experiment, we address the following research question: Can APEX provide high-fidelity simulation for LLM serving? While we demonstrate in Table 2 that APEX can identify optimal parallel execution plans that enhance LLM serving performance, it is also essential to verify the accuracy of its simulation results. We evaluate the simulation fidelity from two perspectives. First, under a fixed number of devices, we assess whether APEX can accurately predict the performance improvement of an optimal plan compared to a baseline plan. Second, for a given execution plan, we evaluate whether APEX can reliably predict performance improvements when scaling across different numbers of devices. We show the experimental results in Figure 5 and Figure 6. We set the arrival rate to 0.5 for both experiments as an example; setting the arrival rate of 0.25 yields similar results.
For the first experiment set, we compare the predicted speedup of Feasible Optimal plans with the actual speedup achieved on the device cluster (Figure 5). We do not evaluate Mixtral-8x22B as vLLM currently does not support expert parallelism. APEX achieves accurate predictions, with an average relative error of 9.5%. The largest discrepancies occur with the Qwen-2.5-32B model, the smallest model in our experiment, where the relative error reaches 23%. This is because, for smaller models, the overhead of operations such as RMSNorm becomes relatively significant, and APEX does not account for these operations in its simulations. In contrast, for larger models like Llama-3.1-70B and Mistral-Large (123B), the predictions are more accurate, as the execution time is dominated by attention and feedforward cells, which are well-modeled by APEX. The fidelity for smaller LLMs could be improved by incorporating profiling results for additional operations. However, as APEX primarily focuses on optimizing performance for serving large models, which are more computationally intensive, we leave this improvement for future work. For the second experiment set, we evaluate the time per output token (TPOT) across different numbers of GPUs using tensor parallelism. We show the results in Figure 6. The APEX prediction results are shown as blue solid lines, while the actual results are represented by red dotted lines. The high similarity between the two lines across all cases demonstrates that APEX accurately captures the scalability trend when scaling from 2 GPUs to 8 GPUs across various models and request traces. The estimated TPOT (ms) of APEX is plotted on the left-hand Y-axis, and the actual TPOT (ms) is plotted on the right-hand Y-axis. The actual measured TPOT is consistently higher than the predicted value, as evident from the larger values on the right-hand Y-axis compared to the left-hand Y-axis. This is because APEX primarily focuses on the overhead of key operations, such as attention and feedforward, while omitting the overhead of other operations. Nevertheless, APEX accurately captures the relative performance differences between various parallel execution plans and degrees of parallelism, as demonstrated in Figure 5 and 6. This capability enables APEX to effectively evaluate and compare different execution plans, thereby identifying an optimal parallel execution plan for LLM serving.
4.4. Evaluation of Efficiency
We evaluate the efficiency of APEX by comparing the actual execution time on the device cluster with the APEX simulation time. Evaluating all parallel execution plans across the setups in Table 2 (i.e., different models, traces, arrival rates) takes approximately 160 hours on 8 H100 GPUs. In contrast, the same evaluation is completed in less than 2.5 hours using APEX on CPU, making it 71$\times$ faster than the actual implementation. From a cost perspective, running the actual implementation would cost approximately \$currency-dollar\\$\$8,889 (based on the Azure NC40ads H100 cluster pricing), whereas the APEX simulation costs only \$currency-dollar\\$\$7.20 (assuming an Azure D64s v6 node). This translates to a 1234.5$\times$ cost reduction, demonstrating the significant efficiency and cost-effectiveness of APEX. Obtaining the operation-level profiling results to set up APEX takes approximately 40 GPU hours. Yet, this is a one-time cost that can be amortized for the same type of hardware device.
Figure 7. APEX maintains similar simulation overhead when scaling from billion-scale to trillion-scale models
We also evaluate the simulation overhead when scaling to larger LLMs on larger device clusters. We use the following example models: Qwen2.5-32B, Llama-3.1-70B, Mistral-Large (123B), and Llama-3.1-405B. We synthesize a trillion-scale model by scaling the Llama-3.1-70B model 16 times; this can be done by modifying values in the LLM configuration file. As shown in Figure 7, APEX demonstrates high scalability, maintaining a similar simulation overhead when scaling from a 32B model to a trillion-scale model. This efficiency is achieved by leveraging the canonical representation in Transformer IR (Section 3.3) and the repetitive structure of the transformer architecture (Section 3.7) to greatly reduce the design space. This scalability highlights APEX’s potential for future applications, as model sizes continue to grow and the cost of actual deployment may become even more expensive.
4.5. Evaluation of Extensibility
LLMs and their serving systems are evolving rapidly. While APEX supports a wide range of models, devices, etc., the simulator may still become outdated as new advancements emerge. To address this, APEX is designed for extensibility, enabling easy adaptation to support the latest developments. We evaluate APEX’s extensibility by measuring the overheads to implement a new feature. We use two metrics: (1) programming overhead, measured in lines of code required for a feature extension, and (2) implementation time overhead, measured in hours, including time to write and execute the necessary code or scripts. We reported the overhead of various types of extensions in Table 3. Below, we provide details of the extensions implemented for the evaluation.
Extending to New Models APEX can effortlessly support a new LLM without additional programming or implementation time by requiring only a configuration file of the model. However, for LLMs containing unknown transformer cells (Section 3.3), such as a novel feedforward network, additional effort is needed. The primary work involves implementing the Parallel Template (Section 3.4) for the new cells. For evaluation, we measured the effort to support SwiGLU cells (Shazeer, 2020) used in the Llama and Qwen models.
Extending to New Device Cluster APEX can easily adapt to new device clusters by providing the device name, memory capacity, and interconnection network (e.g., NVLink, PCIe). We evaluated the effort required to support a new type of GPU and an in-house AI accelerator. The implementation time overhead mainly involves running profiling scripts, which take 4 to 8 hours depending on the hardware’s performance. While extending to new devices requires a relatively long script execution time, this effort is only necessary when a new device is released.
Extending to New Batching Mechanism APEX can also support new batching mechanisms. While this is not as straightforward as supporting a new model or a new device cluster, it can be done by modifying the Iteration-Level Batching Module 3.6. APEX adopts a vLLM-style (Kwon et al., 2023) batching by default. As an example, we evaluate the overhead of extending to support a Sarathi-Serve-style (Agrawal et al., 2024b) batching, which performs chunk prefilling to better interleave prefill and decode requests. This extension is done by adding a new variable, chunk size, to the batching module and a counter to each request to ensure all prefill chunks are completed before moving to the generation stage.
Table 3. Evaluating the overhead of extending APEX
Extension Type Programming Overhead (Lines of Code) Implementation Time Overhead (Hour)
LLM 0 $\sim$0
LLM (w/ unknown cells) 50 - 150 1-2
Device cluster $\sim$20 4-8
Batching Mechansim $\sim$100 1-2
Parallelism 50 - 200 1-2
Extending to New Parallelism The overhead of extending APEX to support new parallelism types depends on the specific parallelism being added. For example, supporting a new type of intra-layer parallelism is analogous to supporting a new transformer cell, as it primarily requires adding a new Parallel Template. Similarly, Fully Sharded Data Parallelism (FSDP) can be integrated into APEX by extending the existing Data Parallelism implementation, incorporating additional collective communication steps, and adjusting memory usage to account for sharded model storage.
In summary, APEX is designed to support a wide range of extensions with minimal programming and implementation overhead, enabling it to stay up-to-date with advancements in LLMs, hardware, and serving systems as they continue to evolve.
4.6. Beyond Identifying Optimal Execution Plan
While APEX is primarily developed to identify optimal parallel execution plans for LLM serving systems, its high simulation fidelity makes it applicable to other use cases as well. Below, we present two examples of applications.
4.6.1. APEX provides insights for building LLM serving platform
When a new hardware device is released, service providers can use APEX to estimate the expected performance improvements from adopting the new hardware by adjusting the hardware parameters of device cluster in APEX. For example, the performance boost of upgrading from Nvidia A100 to H100 GPUs can be projected by scaling the compute time and data transfer time according to their relative compute power and memory bandwidth. Using this approach, APEX estimates that upgrading from A100 to H100 results in a 1.79$\times$ improvement in TTFT and a 1.66$\times$ speedup in TPOT; this estimation closely aligns with actual results reported in (Patel et al., 2024), which shows a 1.85$\times$ improvement in TTFT and a 1.43$\times$ speedup in TPOT. APEX also allows users to freely scale hardware parameters (i.e., creating synthetic hardware) to estimate the hardware upgrade required to achieve a specific performance boost in LLM serving.
4.6.2. APEX provides insights for meeting SLOs
In addition to minimizing end-to-end serving latency, service providers must also meet various service-level objectives (SLOs), such as maintaining a time per output token (TPOT). APEX can assist in achieving these SLOs. A common strategy to meet latency requirements is to adjust the batch size by setting a maximum batch size constraint. APEX begins by simulating a subset of requests to determine an upper bound for the batch size, denoted as $m$. It then divides this upper bound into $n$ segments and simulates the request traces with various maximum batch size constraints, ranging from $\frac{1\times m}{n}$ to $\frac{n\times m}{n}$, where $n$ is determined heuristically. Figure 8 illustrates two examples using the Llama-3.1-70B and Mistral-Large models on the Creation dataset. For simplicity, only results for pure tensor parallelism are shown. However, similar evaluations are performed across all parallel execution plans, resulting in numerous design points to choose from. Assuming a service provider aims to decrease the TPOT to meet the SLO, they can use the results to estimate the necessary adjustment to the maximum batch size constraint. For instance, reducing the maximum batch size constraint from 16 to 8 results in an 18% TPOT improvement for the Llama model and a 14% improvement for the Mistral model. However, overly restricting the maximum batch size can negatively impact end-to-end latency, as illustrated in Figure 8, where the batch size constraint is set to 4.
Figure 8. APEX can simulate with various maximum batch size constraints and suggest adjustments to meet SLOs.
5. Related Work
LLM Serving Systems: With the emergence of LLMs, numerous LLM serving systems have been proposed (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024; Patel et al., 2024). Each system introduces innovations to address key challenges in LLM serving. Orca (Yu et al., 2022) introduced iteration-level batching, significantly improving serving throughput. This technique has since become a standard in LLM serving systems. vLLM (Kwon et al., 2023) proposed PagedAttention, an efficient method to manage KV cache memory usage, enabling more concurrent request batching. Sarathi-Serve (Agrawal et al., 2024b) addressed prefill-decoding interference by proposing chunked prefill, a technique that breaks the prefill stage into smaller stages to enhance batching efficiency; DistServe (Zhong et al., 2024) tackled the same issue by disaggregating the prefill and decoding stages, processing each on separate device clusters optimized for their specific stages. Splitwise (Patel et al., 2024) also disaggregated the two stages, allocating requests to distinct device clusters to achieve higher throughput and cost-effectiveness. While these systems allow users to manually specify parallel execution plans (e.g., degrees of tensor and pipeline parallelism), they do not provide guidance on determining the optimal parallel execution plan. APEX complements these works. Users can utilize APEX to simulate and derive an optimal parallel execution plan tailored to their LLM serving setup, which can then be used to configure the serving system for improved performance.
Partition-Strategy-Search Tools: Several tools have been proposed to identify an optimal configuration for LLM training. Calculon (Isaev et al., 2023) proposes a performance model that helps developers determine an optimal parallel execution plan for LLM training. vTrain (Bang et al., 2024) and ASTRA-sim (Rashidi et al., 2020) are simulators that guide users to find a system configuration that optimizes the LLM training time or the training cost. Nevertheless, these works cannot be applied to LLM serving due to the dynamism of iteration-level batching. LLMServingSim (Cho et al., 2024) and Vidur (Agrawal et al., 2024a) are the few simulation frameworks designed for LLM serving systems. Both perform fine-grained simulations that account for the dynamism of iteration-level batching. However, LLMServingSim primarily focuses on NPUs and Processing-in-Memory (PIM) architectures, as it relies on their corresponding hardware simulators to estimate execution times. In contrast, APEX utilizes operation-level profiling data to estimate execution time and does not rely on other hardware simulators. Vidur requires a model onboarding step before simulation, which involves parsing the operations within the target LLM and profiling them. APEX bypasses this requirement by capturing the key operations of LLMs and utilizing pre-collected profiling results, allowing simulations to start immediately with amortized profiling costs. More importantly, the capabilities of both works are insufficient for simulating state-of-the-art LLM serving systems. Specifically, they do not support quantization techniques, their parallelism support is limited to pipeline and tensor parallelism, and they are constrained to traditional dense LLMs, lacking provisions for emerging architectures such as Mixture of Experts (MoE) models. While it is theoretically possible to extend these simulators to accommodate additional models and parallelism strategies, the required effort remains uncertain due to the lack of evaluations on their extensibility.
6. Conlusion
In this work, we developed APEX, an extensible and dynamism-aware simulator for LLM serving. We evaluated APEX across various LLMs and distinct workloads, demonstrating its high-fidelity simulation with less than 10% relative error on average. APEX successfully identified optimal parallel execution plans, achieving up to 4.42$\times$ speedup compared to heuristic plans. The simulation proved highly efficient, providing a 71$\times$ speedup and 1234$\times$ greater cost-effectiveness than deployment on actual hardware. APEX also showcased high scalability, maintaining a similar simulation overhead when scaling from billion-scale to trillion-scale models. Additionally, we demonstrated APEX’s ease of extensibility to accommodate new models, device clusters, and more. In the future, we plan to extend APEX to support multimodal LLMs, which will involve incorporating parallel execution plans for encoders handling modalities such as vision and audio.

### v2

@@
-Toward High-Performance LLM Serving: A Simulation-Based Approach for Identifying Optimal Parallelism
+APEX: An Extensible and Dynamism-Aware Simulator for Automated Parallel Execution in LLM Serving
 Abstract.
-Large Language Models (LLMs) have achieved significant success in various domains, and serving LLMs efficiently has become crucial.
-LLMs are often served with multiple devices using parallelism techniques like data, pipeline, and tensor parallelisms.
-Each parallelism presents trade-offs between computation, memory, and communication overhead, making it challenging to determine the optimal parallel execution plan.
-Moreover, input workloads also impact parallelism strategies.
-Tasks with long prompts like article summarization are compute-intensive, while tasks with long generation lengths like code generation are often memory-intensive; these differing characteristics result in distinct optimal execution plans.
-Since searching for the optimal plan via actual deployment is prohibitively expensive, we propose APEX, an LLM serving system simulator that efficiently identifies an optimal parallel execution plan.
-APEX performs dynamism-aware simulation, which captures the complex characteristics of iteration-level batching, a technique widely used in state-of-the-art LLM serving systems.
-APEX leverages the repetitive structure of LLMs to reduce design space, maintaining a similar simulation overhead, even when scaling to trillion scale models.
-APEX supports a wide range of LLMs, device clusters, etc., and it can be easily extended to new models or clusters through its high-level templates.
-We run APEX simulations using a CPU and evaluate the optimal parallel execution plans found by APEX using a cluster with 8 H100 GPUs.
-The simulations cover a myriad of LLM architectures, precision formats, cluster configurations, and input workloads.
-We show that APEX can find optimal execution plans that are up to 4.42$\times$ faster than heuristic plans in terms of end-to-end serving latency.
-APEX also reports a set of metrics used in LLM serving systems, such as time per output token (TPOT) and time to first token (TTFT).
-Furthermore, APEX can identify an optimal parallel execution plan within 15 minutes using a CPU.
-This is 71$\times$ faster and 1234$\times$ more cost-effective than actual deployment on a GPU cluster using cloud services.
-APEX will be open-sourced upon acceptance.
-Finally, we show that LLM service providers can utilize APEX to meet service-level objectives and explore hardware design space to build high-performance LLM serving platforms.
+Efficiently serving Large Language Models (LLMs) requires selecting an optimal parallel execution plan, balancing computation, memory, and communication overhead.
+However, determining the best strategy is challenging due to varying parallelism techniques (data, pipeline, tensor) and workload characteristics (e.g., compute-intensive tasks with long prompts vs.
+memory-intensive tasks with long generation).
+We propose APEX, an LLM serving system simulator that efficiently identifies optimal parallel execution plans by considering key factors of LLM serving systems, such as memory usage, batching behavior, etc.
+APEX performs dynamism-aware simulation to model iteration-level batching, and leverages LLMs’ repetitive structure to reduce design space, scaling efficiently to trillion-scale models.
+APEX abstracts the key components of LLM serving systems—including the model, batching module, quantization formats, and device clusters—enabling the simulator to be general and extensible.
+Simulating on a CPU, APEX evaluates execution plans for various device clusters, covering diverse LLMs and workloads.
+APEX finds plans up to 3.37$\times$ faster than heuristics, and also plans that reduce energy consumption by up to 45% compared to latency-optimal plans.
+APEX performs comprehensive evaluations, reporting key system metrics like time per output token and time to first token, which can help service providers meet SLOs.
+APEX identifies an optimal plan within 15 minutes on a CPU, making it 71$\times$ faster and 1234$\times$ more cost-effective than cloud-based GPU deployment.
+APEX can be accessed at: https://github.com/microsoft/apex_plus
 LLM Serving, Optimal Parallelism, Simulation
@@
 Large Language Models (LLMs) have been successfully applied to various applications, such as code generation (Fakhoury et al., 2024; Zhong and Wang, 2024), question-answering (Zhuang et al., 2023; Lucas et al., 2024), and many more (Zhong et al., 2023; Wang et al., 2024; Zhou et al., 2024).
-Serving an LLM is both memory- and compute-intensive; therefore, leveraging a cluster with multiple devices is essential to gain computing power and memory resources for achieving high performance.
-Several techniques have been proposed to parallelize LLM on multiple devices, such as data parallelism (DP) (Li et al., 2020b), pipeline parallelism (PP) (Huang et al., 2019), tensor parallelism (TP) (Shoeybi et al., 2020), etc.
-Each parallelism has its pros and cons.
-Techniques like TP are more memory-efficient than DP as it does not produce model replicas, but they incur expensive collective communication overhead (Shoeybi et al., 2020).
-On the other hand, PP has lower communication overhead but can suffer from workload imbalance (Huang et al., 2019).
-To balance the trade-offs, parallelism techniques can be adopted in a hybrid manner, potentially leading to better serving performance than relying on a single type of parallelism.
-However, determining the optimal parallel execution plan is challenging due to the many factors involved, including computation workload, network traffic, memory efficiency, etc.
+Serving an LLM is both memory- and compute-intensive, necessitating multi-device clusters to achieve high performance.
+Several techniques have been proposed to parallelize LLM, such as data parallelism (DP) (Li et al., 2020b), pipeline parallelism (PP) (Huang et al., 2019), tensor parallelism (TP) (Shoeybi et al., 2020), etc.
+Each parallelism has its pros and cons: TP is memory efficient, as it does not produce model replicas like DP, but it incurs an expensive communication overhead (Shoeybi et al., 2020).
+PP reduces communication overhead, but can suffer from workload imbalance (Huang et al., 2019).
+Hybrid parallelism can effectively balance the trade-offs to improve performance, but identifying the optimal parallel execution plan is non-trivial due to the many factors involved, including computation workload, network traffic, memory efficiency, etc.
 Furthermore, the optimal parallel execution plan also depends on the characteristics of the input requests.
-Some requests lead to long prompts with short generation lengths, such as summarizing an article, while some lead to short prompts and long generation lengths, like code generation and storytelling.
-The former type of request is compute-intensive, while the latter is memory-intensive.
-As a result, these two types of requests favor different parallelism techniques to optimize performance.
+Compute-intensive tasks (e.g., tasks with long prompts like summarization) and memory-intensive tasks (e.g., tasks with long generations like coding) favor different parallelism strategies.
 A common practice is to adopt heuristic execution plans instead, such as applying TP within the same node and PP across different nodes (Miao et al., 2024; vLLM, v060).
-Yet, a recent study has shown that applying such heuristics approaches can be up to 2$\times$ slower than the optimal configuration (Isaev et al., 2023); this calls for the need for service providers to search for and adopt an optimal parallel execution plan rather than relying on heuristics.
-To find an optimal execution plan, a straightforward approach is to deploy and evaluate various parallel execution plans directly; however, this approach is prohibitively expensive, as it could take thousands of GPU hours to assess multiple execution plans (Agrawal et al., 2024b, a).
-Note that this high searching cost cannot be amortized, as the optimal configuration varies depending on the model and the workload characteristics of the input requests.
-An alternative is to search via modeling-based approaches, such as developing performance models (Isaev et al., 2023) or building simulators (Yu et al., 2021; Duan et al., 2024) to estimate the performance of a given parallel execution plan; nevertheless, existing solutions developed for Deep Neural Network (Yu et al., 2021; Duan et al., 2024) or LLM training (Isaev et al., 2023; Bang et al., 2024) cannot be applied to LLM serving, as such systems introduce unique challenges: (1) Dynamism of iteration-level batching: Unlike conventional ML systems that adopt static batching, which waits for all requests in the current batch to be completed before batching new requests, state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) adopt iteration-level batching (Yu et al., 2022) to achieve high serving performance.
-With iteration-level batching, incoming requests are continuously added to the processing batch whenever memory becomes available.
-This makes it challenging to model the system, as the batch size changes dynamically during serving.
-Additionally, some requests within the batch may be in the prefill stage, while others are in the generation stage.
+Yet, recent studies show that heuristics plans can be 2$\times$ slower than optimal configurations (Isaev et al., 2023), underscoring the need for service providers to search for and adopt an optimal parallel execution plan rather than relying on heuristics.
+To find an optimal execution plan, a straightforward way is to deploy and evaluate parallel execution plans exhaustively; however, this is prohibitively expensive as it would take thousands of GPU hours (Agrawal et al., 2024b, a).
+An alternative is to leverage modeling-based approaches, such as performance models (Isaev et al., 2023) or simulators (Yu et al., 2021; Duan et al., 2024), to estimate the performance of parallel execution plans; nevertheless, existing solutions developed for Deep Neural Network (Yu et al., 2021; Duan et al., 2024) or LLM training (Isaev et al., 2023; Bang et al., 2024) cannot be applied to LLM serving, as serving systems introduce unique challenges: (1) Dynamism of iteration-level batching: Unlike conventional ML systems that use static batching, where all requests in a batch must be completed before new ones are added, LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) employ iteration-level batching (Yu et al., 2022).
+In this approach, incoming requests are continuously added to the processing batch as memory becomes available, resulting in dynamically changing batch sizes which are hard to model.
+Moreover, some requests within the batch may be in the prefill stage, while others are in the decode stage.
 This interleaving of stages further complicates the modeling, as these stages have significantly different computational characteristics (details in Section 2.1).
-(2) Exponentially-growing design space: Although modeling-based approaches provide a potentially time-efficient solution to search for the optimal parallel execution plan, they can still incur substantial overhead as the design space grows exponentially with respect to the model size and the number of devices in the cluster.
-Since LLM serving (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) often involves large models and clusters, it is non-trivial to develop a solution that can search for an optimal parallel execution plan within a reasonable timeframe.
+(2) Exponentially-growing design space: The design space grows exponentially with respect to the model size and the number of devices in the cluster.
+Since LLM serving often involves large models and clusters, it is non-trivial to develop a solution that can search for an optimal parallel execution plan within a reasonable timeframe.
 (3) Adapting for continuously-evolving systems: LLM serving system is an emerging area, with continuous evolution in model architecture, hardware platforms, and system optimizations such as quantization (Lin et al., 2024; Hooper et al., 2024) and parallelism techniques (Rajbhandari et al., 2022; Liu et al., 2023).
-Consequently, a performance model or a simulator can easily become obsolete and inapplicable to state-of-the-art LLM serving systems.
+Consequently, a performance model or a simulator can easily become obsolete and inapplicable to the latest LLM serving systems.
 Thus, keeping pace with the rapid evolution of such systems is necessary but also challenging.
 To this end, we propose APEX, an extensible and dynamism-aware simulator for Automated Parallel EXecution in LLM serving.
-APEX takes an LLM, a set of input requests with various context and generation lengths, and a device cluster as inputs, and generates an optimal parallel execution plan for LLM serving.
-Specifically, for a given LLM model and device cluster, APEX first generates various parallel execution plans, each representing a unique way to parallelize the model by combining various parallelism techniques, such as tensor, data, pipeline, and expert parallelisms.
+Given an LLM, a device cluster, and input requests with varying context/generation lengths, APEX generates an optimal parallel execution plan for LLM serving.
+APEX first generates various parallel execution plans, each representing a unique way to parallelize the model by combining various parallelisms (i.e., DP, PP, TP).
 APEX then evaluates each plan by estimating the execution time of serving the input requests through simulation.
-APEX performs dynamism-aware simulation that is capable of modeling the complex characteristics of iteration-level batching, which involves concurrently serving requests of varying lengths and stages (i.e., prefill and generation stage).
-After the simulation, APEX provides a comprehensive evaluation for each parallel execution plan, which includes multiple metrics used in LLM serving systems, such as time per output token (TPOT), time to first token (TTFT), end-to-end serving latency, P95 latency, among others (Miao et al., 2023).
-A parallel execution plan with the lowest end-to-end serving latency is chosen as the optimal plan.
-Despite performing complex dynamism-aware simulations, APEX remains time-efficient by leveraging the repetitive structure of transformer layers in LLMs, which significantly reduces the design space; this allows APEX to scale to trillion-scale models on multi-node device clusters.
-APEX supports a broad range of LLMs, device clusters, quantization formats, and parallelism techniques, which are essential for modeling state-of-the-art LLM serving systems.
-Recognizing the rapid evolution of LLM serving systems, APEX is designed to be modular and extensible.
-We capture the high-level abstractions of LLMs and device clusters, and develop software templates based on these abstractions.
-This approach allows new models and device clusters to be easily supported with minimal coding effort in the templates.
-Additionally, APEX leverages operation-level profiling results to estimate the execution time of parallel execution plans on a given device cluster.
-Profiling-based estimation enables APEX to adapt to different clusters by collecting profiling data from the underlying platform.
-Note that the device cluster APEX supports is not limited to GPU clusters, but also AI-accelerator clusters such as TPU (Jouppi et al., 2023), Intel Gaudi (Intel, 2024), etc.
-We plan to open-source APEX upon paper acceptance.
-We further demonstrate that, in addition to identifying optimal parallel execution plans, APEX can also assist LLM service providers in meeting service-level objectives (SLOs) and provide valuable insights for building LLM-serving hardware platforms.
-The key contributions of this work are:
-  • We propose APEX, a dynamism-aware simulator that captures the complex characteristics of LLM serving.
-APEX automatically finds an optimal parallel execution plan for a given LLM, a device cluster, and a set of input requests.
-  • APEX supports a wide range of LLMs, parallelism techniques, quantization formats, and device clusters. APEX can also be easily extended to new models and device clusters.
-  • We evaluate APEX using four LLMs on three datasets with distinct workloads.
-APEX identifies optimal parallel execution plans that are up to 4.42$\times$ faster than heuristic plans in terms of the end-to-end serving latency.
-  • APEX provides high-fidelity simulation, achieving less than 10% relative error on average when predicting the speedup of optimal execution plans over the baseline plans.
-  • APEX is time-efficient and cost-effective, capable of finding an optimal parallel execution plan within 15 minutes on a CPU; this is 71$\times$ faster and 1234$\times$ cost-effective than actual deployment on a GPU cluster using cloud services.
-  • APEX is highly scalable, maintaining similar simulation overhead when scaling from billion-scale to trillion-scale models.
-  • APEX features comprehensive evaluation, which can help service providers meet SLOs. APEX also provides insightful suggestions for building LLM-serving hardware platforms.
+APEX performs dynamism-aware simulation that is capable of modeling the characteristics of iteration-level batching.
+After simulation, APEX provides a comprehensive evaluation for each execution plan with key LLM serving metrics, such as time per output token (TPOT), time to first token (TTFT), and energy consumption, among others (Miao et al., 2023).
+This detailed evaluation enables service providers to select execution plans optimized for various objectives (e.g., minimizing latency or energy consumption) while meeting service-level objectives (SLOs).
+Despite performing complex simulations, APEX remains highly scalable by leveraging the repetitive structure of transformer models, significantly reducing design space and enabling simulations at trillion-scale models across multi-node clusters.
+To model state-of-the-art LLM serving systems, APEX supports a broad range of LLMs, parallelisms, quantization formats, and device clusters.
+Recognizing the rapid evolution of LLM serving, APEX is designed to be modular and extensible, allowing easy integration of new models and hardware with minimal effort.
+Additionally, APEX leverages operation-level profilings to estimate the system execution time, which allows APEX to adapt to different clusters by collecting performance data from the underlying platform.
+Note that APEX not only supports GPU clusters but also ASIC clusters such as TPU (Jouppi et al., 2023) and Intel Gaudi (Intel, 2024).
+We evaluate APEX using a myriad of LLMs, device clusters, and distinct workloads.
+APEX consistently identifies optimal execution plans that outperform heuristic plans while maintaining low simulation overhead.
+APEX is open-sourced: https://github.com/microsoft/apex_plus.
+Our key contributions are:
+  • Automated parallel execution: We propose APEX, a simulator that automatically finds an optimal parallel execution plan for a given LLM, a device cluster, and input requests.
+  • Dynamism-aware simulation: APEX models iteration-level batching, capturing real-world LLM serving behavior.
+  • Accurate & effective: APEX delivers accurate speedup predictions, with an average relative error of just 10.7%.
+APEX finds optimal plans up to 3.37$\times$ faster than heuristic plans, and energy-optimal plans that reduce energy consumption by up to 45% compared to latency-optimal plans.
+  • Broad compatibility and extensibility: APEX supports various LLMs, parallelisms, quantizations, and device clusters, while allowing easy extension to new configurations.
+  • Cost-effective & scalable: APEX identifies an optimal plan within 15 minutes on a CPU, making it 71$\times$ faster and 1234$\times$ more cost-effective than cloud-based GPU deployment.
+It also maintains similar simulation overhead when scaling from billion-scale to trillion-scale models.
+  • Comprehensive evaluation: APEX reports key LLM-serving metrics, which can help service providers meet SLOs.
 2. Background
 2.1. LLM Inference
-Large Language Models (LLMs) are a type of ML models that are built upon transformer architecture (Vaswani et al., 2017).
-LLM inference takes a user prompt as input, and generates a response, token by token, sequentially.
-The sequential token generation process is also known as the auto-regressive generation, which utilizes the previously generated tokens as input to predict the next token.
+Large Language Model (LLM) inference takes a user prompt as input, and generates a response token by token.
+The sequential token generation process is also known as auto-regressive generation (Gregor et al., 2014), which utilizes previously generated tokens as input to predict the next token.
 The inference process of LLMs consists of two stages:
+1.
 Prefill Stage: In the prefill stage, LLM processes the input request (i.e., prompt) to set up intermediate states (keys and values) that are used to predict the first token.
@@
 This high degree of parallelism makes the prefill stage compute-bound (Zhong et al., 2024).
-Generation Stage: During the generation phase, LLM generates output tokens autoregressively until a stopping criterion is met.
+2.
+Decode Stage: During the decode stage, LLM generates output tokens autoregressively until a stopping criterion is met.
 The generation of each token depends on all previous tokens’ output states (keys and values).
-The generation phase is often memory-bound (Lin et al., 2024), as the latency mainly depends on the speed of data transfer for the output states from the memory, rather than on the computation itself.
-2.2. LLM Serving and Iteration-Level Batching
-LLMs are often hosted by service providers in the cloud (Li et al., 2024).
-Users send requests to these providers through APIs or chatbots and then receive responses generated by the LLMs.
-While a single LLM inference processes a single input request and generates one output response, LLM serving concurrently processes multiple input requests from the users and generates multiple output responses, i.e., multiple LLM inferences happen in parallel.
-For LLM serving, it is critical to achieve both low latency and high throughput: low latency is necessary to meet service-level objectives, while high throughput enables service providers to serve a large number of users simultaneously.
-Due to the autoregressive nature of LLMs, the lengths of the generated responses can vary significantly.
-Consequently, the commonly used static batching leads to suboptimal performance in LLM serving, as all the batched requests need to wait for the longest response to be generated to proceed to process new requests.
-To overcome this inefficiency, iteration-level batching (Yu et al., 2022) is proposed.
-Iteration-level batching continuously schedules newly arrived request(s) into the existing batch whenever GPU memory becomes available, rather than waiting for all the requests to be completed.
-Iteration-level batching significantly improves the serving throughput and is widely adopted in state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024).
+The generation phase is often memory-bound (Lin et al., 2024), as its latency is primarily determined by the data transfer time of reading the output states from memory.
 2.2. Device Clusters and Interconnections
-Serving an LLM is both memory- and compute-intensive.
 To achieve high serving throughput and accommodate the large model size, LLMs are often deployed on a cluster of devices.
@@
 Devices within the same node are typically connected by high-bandwidth interconnects like NVLink (Li et al., 2020a) at level 1, while devices across different nodes are connected via inter-node networks such as InfiniBand (Liu and Guitart, 2022) at level 2.
-In addition to GPUs, clusters of AI accelerators are also emerging, such as TPU clusters (Jouppi et al., 2023) and Gaudi clusters (Intel, 2024).
+In addition to GPUs, ASIC clusters are also emerging, such as TPU and Gaudi clusters (Jouppi et al., 2023; Intel, 2024).
 These clusters also utilize tree-based network topologies for interconnection and can be abstracted similarly as GPU clusters (Choquette, 2023).
-Figure 1. An example of a two-level device cluster hierarchy, where leaves represent the devices. The memory bandwidth and latency are uniform within the same level.
-Figure 2. System overview of APEX
+2.3. LLM Serving and Iteration-Level Batching
+LLMs are often hosted by service providers in the cloud (Li et al., 2024).
+Users send requests to these providers through APIs or chatbots and then receive responses generated by the LLMs.
+For LLM serving, it is critical to achieve both low latency and high throughput: low latency is necessary to meet service-level objectives, while high throughput enables service providers to serve a large number of users simultaneously.
+Due to the auto-regressive nature of LLMs, the lengths of the generated responses can vary significantly.
+Consequently, the commonly used static batching leads to suboptimal performance in LLM serving, as all the batched requests need to wait for the longest response to be generated to proceed to batch new requests.
+To overcome this inefficiency, iteration-level batching (Yu et al., 2022) is proposed.
+Iteration-level batching continuously schedules newly arrived request(s) into the existing batch whenever GPU memory becomes available, rather than waiting for all the requests to be completed.
+Iteration-level batching significantly improves the serving throughput and is widely adopted in state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024).
+Figure 1. An example two-level device cluster. Memory bandwidth and latency are uniform within the same level.
 2.4. LLM Parallelisms
 LLM is often served with multiple devices (Figure 1), and various approaches have been proposed to parallelize LLM on the cluster. We discuss several representative and widely-used parallelisms.
-Data Parallelism (DP): In DP (Li et al., 2020b), the model is replicated across multiple devices.
-The input requests are split into micro-batches and distributed to each model replica for processing.
-DP incurs no communication overhead as each micro-batch is processed independently.
-However, having model replicas incurs large memory overhead, and reduces the number of requests that can be batched concurrently.
+Data Parallelism (DP): In DP (Li et al., 2020b), the model is replicated across multiple devices, and input requests are divided into mini-batches, each processed independently by a replica.
+This approach avoids communication overhead during execution, but comes at the cost of increased memory usage due to model replication, thereby reducing available memory for batching requests.
 Pipeline Parallelism (PP): PP (Huang et al., 2019; Narayanan et al., 2019) divides the model in a layer-wise fashion.
@@
 Yet, TP incurs high collective communication overhead, such as AllReduce, which is prohibitively expensive for interconnection networks with limited bandwidth like PCIe.
-Expert Parallelism (EP): EP (Rajbhandari et al., 2022) is a special type of parallelism for Mixture of Experts (MoE) models.
+Expert Parallelism (EP): EP (Rajbhandari et al., 2022) is a special type of parallelism for Mixture of Experts (MoE) models (Cai et al., 2024).
 EP deploys different experts across various devices.
@@
 3. APEX Simulator Design
+Figure 2. System overview of APEX
 3.1. System Overview
-We depict the system overview of APEX in Figure 2.
-Initially, the Offline Profiler (Section 3.2) of APEX takes a set of profiling scripts to obtain the performance information of the underlying platform, and the results are stored for APEX’s profiling-based simulation.
-Profiling is an offline process that is only performed once when porting to a new cluster with unknown devices (e.g., porting from an A100 to an H100 GPU cluster.) Next, given an LLM and device cluster, the Parallel Execution Plan Generator converts the model into an intermediate representation (Section 3.3) and generates various parallel execution plans (Section 3.4 and 3.5); each plan represents a unique way to map the LLM onto the device cluster.
-The Iteration-Level Batching Module (Section 3.6) takes the generated parallel execution plans and a set of request traces as inputs, and starts batching the input requests.
-The Batching Module determines whether a request should be added to or removed from the batch in each iteration, and reports the active requests to the Dynamism-Aware Simulator (Section 3.7).
-Based on the active requests in the batch, the Dynamism-Aware Simulator estimates the execution time for each iteration and keeps track of multiple LLM system metrics such as time to first token (TTFT), time per output token (TPOT), P95 latency, etc.
-After all requests have been processed, the Dynamism-Aware Simulator produces a Simulation Report for the corresponding parallel execution plan.
-This process is repeated for all the generated parallel execution plans, and the plan with the lowest end-to-end latency is selected as the optimal plan.
-3.2. Offline Profiler
-The Offline Profiler collects performance information of the underlying platform, where the results are used by the Dynamism-Aware Simulator (Section 3.7) to estimate the execution time of different parallel execution plans.
-The Offline Profiler leverages the fact that LLMs are based on transformer architecture, which consists of similar operations, and performs operation-level profiling, such as measuring the computation time of multi-head attention, which is used in the attention layers, and general matrix multiplication (GEMM), which is used in the feedforward layers.
-Profiling the key transformer operations allows APEX to support various LLMs without exhaustively profiling each model, as they can be broken down into the same set of operations.
-In addition to computation operation, the Offline Profiler also profiles the collective communication overheads on the cluster, e.g., the time to perform AllReduce, ReduceScatter, All-to-All.
-This allows APEX to support various parallelisms, which require different collective communications.
-For the computation operations, the profiling is performed by going through various sequence lengths, number of attention heads, hidden dimensions, etc.
-For the collective communication operations, the profiling is performed by varying the data transfer sizes and also the number of devices involved, both within a node and across nodes, in the collective operation.
-Operation-level profiling produces profiling results that can support various LLMs, request traces, parallelisms, and device clusters of different sizes.
-The profiling only needs to be performed once when porting to a new cluster with unknown devices, such as porting from an A100 GPU cluster to a H100 GPU cluster.
-Therefore, for the same type of device cluster, the profiling overhead is a one-time cost that can be amortized.
+We depict the workflow of APEX in Figure 2.
+User-provided input data is shown in yellow-colored boxes, and built-in system data is represented by red-colored cylinders.
+Given an LLM and a device cluster, APEX generates multiple parallel execution plans using built-in Parallel Templates.
+Each generated executon plan represents a distinct strategy for mapping the LLM across the cluster.
+The Batching Module takes these generated plans and a trace of requests as inputs, dynamically generating the processing batch based on request arrival time and resource constraints.
+Specifically, the Batching Module determines, at each iteration, whether a request should be added to or removed from the processing batch and reports the active requests to the LLM Serving Simulator; the Simulator then uses profiling results to estimate execution time and energy consumption for processing the active requests, while also tracking system metrics such as time to first token (TTFT), time per output token (TPOT), P95 latency, among others.
+Once all requests are processed, the simulator generates a simulation report for each parallel execution plan.
+APEX can optimize towards different objectives by selecting the most suitable execution plan based on a parametrizable target metric: It can optimize for energy-efficiency by minimizing energy per inference, or can prioritize throughput by reducing end-to-end latency, enabling flexible adaptation to diverse system goals.
+Suppose that we aim to find the latency-optimal plan for serving Llama-3.1 70B on a cluster with 8 Nvidia H100 GPUs.
+APEX first generates multiple execution plans, such as 8-way tensor parallelism, or a 4-way tensor parallelism combined with 2-way pipeline parallelism.
+Next, it utilizes a trace of requests as input for simulation, such as the Azure LLM Traces (Patel et al., 2024).
+APEX simulates iteration-level batching to organize arriving requests into the processing batch.
+Finally, the LLM Serving Simulator estimates the execution time of each plan and select the one with the lowest end-to-end latency for serving Llama-3.1 70B on the cluster.
+Figure 3. The Parallel Execution Plan Generator produces plans that map a given LLM onto the target cluster using distinct parallelization strategies
+3.2. Parallel Execution Plan Generator
+In this section, we describe the details of the Parallel Execution Plan Generator, which constructs various parallel execution plans to map the input LLM onto the target device cluster.
+The workflow of the Plan Generator is illustrated in Figure 3.
+The process begins by converting the input LLM into Transformer IR, an intermediate representation that provides a canonical abstraction for transformer-based models.
+Using this representation, the Plan Generator produces multiple parallel schemes based on the Parallel Templates.
+A parallel scheme maps the model onto a logical device cluster (Zhao et al., 2020), a virtual overlay that abstracts hardware details without network topology information.
+Next, the Device Mapper assigns logical devices to physical devices, generating executable parallel execution plans, which serve as inputs for the Batching Module.
+This two-stage mapping approach (first to the logical and then to physical device cluster) reduces the design complexity.
+We discuss each module in detail below.
+3.2.1. IR Converter and Transformer IR
+The Plan Generator converts an LLM into Transformer IR, a canonical representation of transformer-based models, as illustrated in Figure 4.
+We develop an IR converter that parses an LLM’s configuration file, extracting parameters such as the number of attention heads and layers, and translates the model into Transformer IR.
+In Transformer IR, an LLM is represented as a chain of cells, where each cell corresponds to key transformer operations, such as multi-head attention or the feedforward network.
+For example, GPT-3 (Brown et al., 2020) can be represented as a sequence of multi-head attention (MHA) cells and multi-layer perceptron (MLP) cells, while Llama-3.1 (Dubey et al., 2024) is modeled as a chain of group query attention (GQA) cells (Ainslie et al., 2023) and SwiGLU cells (Shazeer, 2020).
+Each cell consists of multiple tasks, such as individual attention heads in an MHA cell or experts in a mixture-of-experts (MoE) cell.
+Given that LLMs typically consist of repeated structures, we define the smallest set of nonrepetitive adjacent cells as a block.
+Using Transformer IR, an LLM is uniformly represented as multiple identical block, where each block consists of cells, and each cell contains multiple tasks.
+Transformer IR abstracts an LLM by focusing solely on key operations like MHA and MLP, while omitting operations like tokenization and position embeddings, which are less relevant for model parallelization.
+By focusing on key operations and leveraging the repetitive nature of LLMs through blocks, Transformer IR reduces both the search space and the simulation runtime for APEX.
+Furthermore, Transformer IR serves as a foundation for developing our Parallel Templates, which we describe in the following subsection.
 Figure 4. Transformer IR represents LLMs in a canonical way
-3.3. Transformer IR
-APEX identifies an optimal parallel execution plan by generating and evaluating multiple candidate execution plans.
-Yet, it is non-trivial to generate parallel execution plans for a variety of LLMs as they have different model architectures.
-To address this, we introduce the Transformer IR, a unified abstraction for transformer-based models.
-APEX utilizes Parallel Templates, developed based on the Transformer IR rather than specific models, to generate parallel execution plans.
-This approach enables APEX to parallelize a wide range of LLMs that can be represented through the Transformer IR.
-We discuss the details of Parallel Templates in Section 3.4.
-We depict the idea of Transformer IR in Figure 3.
-We define the key operations of transformers, such as the multi-head attention, as cells.
-An LLM can be represented as a chain of cells.
-For example, GPT-3 (Brown et al., 2020) models can be represented as a chain of multi-head attention cells and multi-layer perceptron cells, and Llama-3.1 (Dubey et al., 2024) models can be represented as a chain of group query attention (Ainslie et al., 2023) cells and SwiGLU (Shazeer, 2020) cells.
-The Transformer IR represents an LLM canonically, which only captures the key cells in the model and ignores operations like tokenization and position embeddings, as they are less relevant for model parallelization; this reduces the search space for simulation.
-Given that LLMs often use the same set of cells repeatedly, we further define the smallest set of non-repetitive adjacent cells as a block.
-Each cell consists of multiple tasks, such as the attention heads in the multi-head attention cell, or the experts in the mixture-of-expert cell.
-Each task works independently from the other tasks (i.e., no inter-communication is needed), and their outputs are combined via operations like concatenation or AllReduce.
-To summarize, using the Transformer IR, an LLM can be unifiedly represented as multiple identical blocks, each block consists of numerous cells, and each cell consists of multiple tasks.
-We develop an IR converter, which parses the information of an LLM configuration file, such as the number of attention heads, number of layers, etc., and represents the model using Transformer IR.
-3.4. Parallel Templates
-We develop Parallel Templates to generate various parallel execution plans for LLM serving.
-As mentioned in Section 3.3, the templates are developed based on the Transformer IR, which allows the templates to support a wide range of LLMs.
-Each cell type is associated with a pre-defined Parallel Template that describes how it can be parallelized on multiple devices.
-We show some examples in Figure 4.
-If tensor parallelism is applied to both the multi-head attention (MHA) cell and the multi-layer perceptron (MLP) cell, as in Figure 4 (a), the tasks (e.g., attention heads) in the two cells are evenly distributed to each device.
-In addition to distributing the tasks, the Parallel Templates also handle the collective communications between adjacent cells.
-In the example of Figure 4 (a), an AllReduce is performed between the cells for data synchronization, which is similar to the case in Megatron-LM (Shoeybi et al., 2020).
-If the numbers of cell replicas vary between the adjacent cells, tensor resharding is required for the output activation; such cases happen when different data parallelism (DP) degrees are applied.
-For example, in Figure 4 (b), a 2-way DP is applied to the MLP cell, and a 1-way DP (i.e., no DP) is applied to the MHA cell, resulting in a different number of cells replicas.
+3.2.2. Parallel Templates and Parallel Schemes Generator
+We develop Parallel Templates to generate parallel execution plans for LLM serving.
+Given the wide variety of LLMs, exhaustively designing templates for each model is impractical.
+Instead, we develop our Parallel Templates based on the Transformer IR, which provides a unified abstraction of LLMs.
+This approach enables the developed templates to support a broad range of models rather than being tailored to specific ones.
+Figure 5 illustrates examples of our templates.
+Each cell type is associated with a pre-defined template that specifies how it is parallelized across multiple devices.
+For instance, if tensor parallelism (TP) is applied to both the multi-head attention (MHA) cell and the multi-layer perceptron (MLP) cell, as shown in Figure 5 (a), tasks (e.g., attention heads) are evenly distributed across devices.
+Beyond task distribution, Parallel Templates also manage collective communication between adjacent cells.
+In Figure 5 (a), an AllReduce operation synchronizes data between cells, similar to the approach in Megatron-LM (Shoeybi et al., 2020).
+In Figure 5 (b), 2-way data parallelism (DP) is applied to the MLP cell, while no DP is applied to the MHA cell, leading to distinct model partitioning.
 Consequently, All-to-All and AllGather operations are performed for tensor resharding.
-Figure 4 (c) depicts a Parallel Template associated with the MHA cell: $H$ attention heads are evenly distributed to $D$ device for computation, and collective communications for resharding and synchronization may also be performed, depending on the number of cell replicas of the adjacent cells.
-Note that while Figure 4 shows a simplified example of having only two devices, the Parallel Templates are parameterized and can support any number of devices.
+While Figures 5 (a) and (b) depict simplified examples using two devices, Parallel Templates are fully parameterized to support any number of devices, as in Figure 5 (c), where $H$ attention heads are evenly distributed across $D$ devices.
 Figure 5. Examples of APEX’s Parallel Templates for two devices. The parameterized templates can extend to D devices.
-3.5. Parallel Scheme Generator & Device Mapper
-Given an LLM and pre-defined Parallel Templates, the Parallel Scheme Generator produces various parallel schemes, which maps the LLM onto a logical device cluster (Zhao et al., 2020) with various combinations of parallelism techniques.
-A logical device cluster is a virtual overlay that describes the number of devices in the cluster without the information of the network topology.
-Consequently, a parallel scheme only defines the parallelisms applied to the Transformer IR cells and their corresponding collective communications, such as the examples in Figure 4 (a) and (b), without assigning the cells to specific physical devices.
-We introduce a virtual overlay of a logical device cluster for design simplicity; this approach allows us to decouple the Parallel Template designs (Section 3.4) from the hierarchal network topology of a device cluster - jointly considering the network topology in the Parallel Templates would greatly increase design complexity.
-The Parallel Scheme Generator creates parallel schemes using a hierarchical top-down approach, beginning with the most coarse-grained model-level parallelism and gradually refining to fine-grained cell-level parallelism.
-The detailed procedure is outlined in Algorithm 1.
-First, the Generator decides the degree of model-level data parallelism (i.e., the number of model replicas), and also derives the number of devices assigned to each replica (i.e., number of replica devices).
-APEX requires the device cluster to be evenly partitioned by the number of model replicas, so the available degrees of parallelism are restricted to the divisors of the total number of devices in the cluster.
-Second, the Generator decides the number of pipeline stages within each replica.
-Similarly, APEX requires the cluster to be evenly partitioned by the number of stages, so the available parallelism degrees are the divisors of the number of devices assigned to the replica.
-Third, the Generator decides the degree of cell-level data parallelism for each cell, and generates task mappings using the pre-defined Parallel Templates.
-If the degree of cell-level data parallelism for a specific cell is less than the number of assigned devices, the Parallel Templates apply intra-layer parallelism like tensor or expert parallelism to parallelize the cell (Section 3.4).
-For example, if four devices are assigned to a cell with two replicas, two devices are assigned to each replica, and the cell can be parallelized using tensor parallelism.
-Afterward, for each pair of adjacent cells, the Generator inserts collective communications into the parallel schemes if tensor resharding is needed.
+Given Parallel Templates and an LLM represented in Transformer IR, the Parallel Scheme Generator constructs parallel schemes that map the LLM onto a logical device cluster (Zhao et al., 2020) using different combinations of parallelism techniques.
+A logical device cluster is a virtual overlay that specifies the number of devices in a cluster without considering network topology.
+As a result, a parallel scheme defines the parallelism applied to Transformer IR cells and their collective communications (e.g., Figure 5 (a), (b)) without assigning the cells to specific physical devices.
+Introducing logical device clusters simplifies design complexity by decoupling Parallel Templates from the hierarchical network topology of physical clusters; the Device Mapper later handles physical mappings.
+The Parallel Scheme Generator follows a hierarchical top-down approach, starting with coarse-grained model-level parallelism and progressively refining to fine-grained cell-level parallelism.
+As detailed in Algorithm 1, the Generator first selects the degree of model-level data parallelism (i.e., number of model replicas) and partitions the cluster accordingly.
+APEX enforces even partitioning, restricting parallelism choices to divisors of the total number of devices.
+Next, the Generator determines the number of pipeline stages within each replica, again ensuring even partitioning.
+Finally, it assigns cell-level data parallelism and applies Parallel Templates to generate task mappings.
+If a cell’s data parallelism degree is lower than the number of assigned devices, intra-layer parallelism (e.g., tensor or expert parallelism) is applied.
+For instance, if four devices are assigned to a cell with two replicas, each replica receives two devices, and tensor parallelism is applied to parallelize the cell replica across the two devices.
+Lastly, for each pair of adjacent cells, the Generator inserts collective communication operations for synchronization and tensor resharding.
+Now that various parallel schemes have been generated for the target model, the logical devices can be mapped to physical devices, producing executable parallel execution plans; we discuss the physical device mapping in the following subsection.
 Algorithm 1 Parallel Scheme Generation
@@
 16:         parallel_schemes.append(stage_schemes)
-After generating the parallel schemes, the Device Mapper maps the logical devices onto the physical device cluster, producing parallel execution plans.
+3.2.3. Device Mapper
+The Device Mapper takes the parallel schemes and the device cluster configuration as input, and assigns logical devices cluster to the physical device cluster, producing parallel execution plans.
 While the parallel schemes are generated in a top-down approach, the Device Mapper operates in a bottom-up manner.
-The logical devices are first mapped to physical devices connected at the bottom level of the cluster (see Figure 1); if the number of logical devices exceeds the available physical devices at the current level, the Device Mapper moves to the next upper level to include additional physical devices for mapping.
-Since lower-level device connections generally offer higher bandwidth than upper-level connections, the Device Mapper prioritizes mapping logical devices assigned to the same cell, as fine-grained cell-level parallelisms tend to incur expensive collective communication overhead, such as the AllReduce in tensor parallelism.
-The Device Mapper then maps logical devices assigned to the same pipeline stage, followed by those assigned to the same model replica, progressing through increasingly coarse-grained parallelism levels.
-This bottom-up approach maps logical devices with potentially higher communication overhead to physical devices connected at the lower level, which have higher interconnection bandwidth, and vice versa.
-3.6. Iteration-Level Batching Module
-The Iteration-Level Batching Module takes the candidate execution plans and request traces as inputs, and simulates the behavior of a serving system using iteration-level batching (Yu et al., 2022).
-Each request comprises three key attributes: context length, generation length, and arrival time.
-As mentioned in Section 2.2, iteration-level batching continuously adds arriving requests to the current batch whenever memory permits.
-Therefore, a Resource Monitor keeps track of the memory usage and determines whether a new request can be batched.
-The Request Scheduler also checks the arrival time of a request to determine whether a request has arrived and can be batched.
-The Request Scheduler maintains a list of active requests, which are the requests in the current batch.
-The active request list also tracks how many tokens have been generated for each request.
-If the number of generated tokens matches the generation length of a request, then the request is completed and should be removed from the active batch; this releases memory to batch new requests.
-One token is generated for each active request in one iteration.
-In each iteration, the Iteration-Level Batching Module sends the active request list to the Dynamism-Aware Simulator (Section 3.7) to estimate the execution time.
-The Request Scheduler greedily adds arriving requests to the batch whenever memory is available to accommodate the context length of a request, without considering the memory required for the KV cache of generated tokens.
-As a result, memory capacity may be exhausted before token generation for all requests is completed.
-In such cases, the most recently added requests and their generated tokens are temporarily removed from the batch to free up memory for earlier requests to complete their token generation.
-These removed requests are re-added to the batch as memory becomes available.
-Both the removal and re-addition are performed in order, prioritizing completing the earliest requests.
-3.7. Dynamism-Aware Simulator
-Given the active requests produced by the Iteration-Level Batching Module and the Profiling Results, the Dynamism-Aware Simulator estimates the execution time of each iteration.
-The active request list contains information such as the context length of the request and the number of tokens that have already been generated for a request.
-This allows the Simulator to identify whether a request is in the prefill stage or the generation stage.
-Specifically, a request is in the prefill stage if the number of generated tokens equals zero; otherwise, it is in the generation stage.
-For requests in the prefill stage, we calculate the execution time of the operations by setting the input sequence length as the context length, as the entire sequence is processed in parallel (Section 2.1).
-For requests in the generation stage, we calculate the execution time of the operations by setting the input sequence length as one; this is because the output states (i.e., key and value) of the previous tokens are stored in the KV cache and do not require recomputation.
-Thus, the computation is only performed on the current token, reducing the effective sequence length to one.
-Furthermore, requests in the generation stage are processed in parallel; thus, assuming $n$ generation requests, the Simulator estimates the execution time by setting the sequence length as $n$, as each request has an effective sequence length of one.
-The Profiling Results contain the execution time of different attention heads, hidden dimensions, sequence lengths, etc., allowing the Simulator to estimate the total execution time for processing the prefill requests and the generation requests.
-If a specific data point is not presented in the Profiling Results, the Simulator estimates its value by leveraging linear interpolation between the nearest available profiling points.
-The Simulator is dynamism-aware as the Iteration-Level Batching Module continuously updates the list of active requests in every iteration.
-In addition, the Simulator utilizes the repetitive nature of LLMs to reduce simulation overhead.
-As discussed in Section 3.3, a block in Transformer IR is the smallest set of non-repetitive cells.
-Thus, the Simulator performs simulation using a single block and projects the execution time of the entire model, which consists of multiple repetitive blocks.
-The projection calculation, such as summing or taking the maximum of the execution times, depends on the connection pattern of the blocks (e.g., sequential or pipelined).
-In addition to estimating the execution time of each iteration, the Simulator also features a Metric Monitor, which calculates the value of several important metrics in LLM serving systems, such as time to first token (TTFT), time per output token (TPOT), P95 latency, Model FLOPs Utilization (MFU), Model Bandwidth Utilization (MBU), among others.
+Specifically, the logical devices are first mapped to physical devices connected at the bottom level of the cluster, such as Level 1 in Figure 1; if the number of logical devices exceeds the available physical devices at the current level, the mapping extends to the next higher level to include additional physical devices.
+Since lower-level connections generally offer higher bandwidth (e.g., NVLink in Level 1), the Device Mapper prioritizes mapping logical devices assigned to the same cell, as fine-grained parallelism (e.g., tensor parallelism) often incurs expensive collective communication (e.g., AllReduce).
+Next, the Device Mapper maps logical devices assigned to the same pipeline stage, followed by those assigned to the same model replica, progressing through increasingly coarse-grained parallelisms.
+This bottom-up approach ensures that logical devices with higher communication demands are mapped to lower-level physical connections with greater bandwidth, while those with lower communication overhead are placed at higher levels.
+Once the mapping is completed, we obtain multiple parallel execution plans, each describing a distinct strategy for parallelizing the input LLM on the target cluster.
+Next, these plans are evaluated using the Batching Module and the LLM Serving Simulator.
+3.3. Batching Module
+The Batching Module takes the parallel execution plans and a trace of requests as inputs, and simulates LLM serving system using iteration-level batching (Yu et al., 2022).
+Request traces are often collected from online services like Azure LLM Inference traces (Patel et al., 2024), and each request comprises three key attributes: context length, generation length, and arrival time.
+As mentioned in Section 2.3, iteration-level batching greedily adds arriving requests to the processing batch whenever memory permits.
+Therefore, the Batching Module tracks memory usage and request arrival times to determine whether new requests can be added to the processing batch.
+It maintains an active request list that tracks requests that are in the processing batch and the length of their generated tokens.
+When a request reaches its generation length, it is removed from the batch, freeing memory for new requests.
+In each iteration, one token is generated per active request, and the active request list is sent to the LLM Serving Simulator (Section 3.4) to estimate the execution time and energy consumption of that iteration.
+Due to the greedy nature of iteration-level batching, the Batching Module may add requests that exceed memory capacity, as it does not pre-allocate enough memory for the KV cache of generated tokens.
+In such cases, the most recently added requests and their tokens are temporarily removed to free memory for earlier requests to complete.
+3.4. LLM Serving Simulator
+The LLM Serving Simulator estimates the execution time and energy consumption for each iteration based on active requests from the Batching Module and Profiling Results.
+The Profiling Results provide execution times and energy consumptions for key operations such as multi-head attention under various configurations (e.g., number of attention heads, feature dimensions).
+If a specific data point is missing, the Simulator applies linear interpolation between the nearest profiling data points.
+We discuss the profiling process in detail in the next subsection and focus here on the Simulator’s execution process.
+The Simulator estimates the time to process the requests in the active request list.
+For requests in prefill stage, execution time is estimated by querying the profiling results with the full context length, as the entire sequence is processed in parallel.
+For requests in decode stage, only the current token is processed, reducing the effective context length to one.
+If there are $n$ decode requests, execution time is estimated using context length $n$, since all decode requests (each with an effective context length of one) are processed in parallel.
+The Simulator estimates energy consumption in a similar fashion by querying the profiling results for the energy consumed by each operation.
+To reduce simulation overhead, the Simulator leverages the repetitive structure of LLMs by simulating only a single Transformer block, which is the smallest set of non-repetitive cells of a LLM (Section 3.2.1).
+APEX then extrapolates the results to estimate the execution time and energy consumption of the full model.
+For example, the execution time of the iteration is extrapolated by taking the maximum across all pipeline stages, since the latency is determined by the slowest pipeline stage.
+In contrast, energy consumption is computed as the sum across all stages, as all devices contribute to energy usage.
+The Simulator also tracks key LLM serving metrics such as time to first token, time per output token, P95 latency, Model FLOPs Utilization, and Model Bandwidth Utilization; these metrics allow APEX to evaluate the parallel execution plans comprehensively and pick an optimal plan based on a user’s desired metric(s).
+3.5. Profiling Results
+The profiling results are collected offline and used as system data (Figure 2) during simulation.
+An Offline Profiler collects platform-specific performance data, which is used by the LLM Serving Simulator to estimate execution time and energy consumption.
+Instead of profiling each LLM exhaustively, the Profiler performs operation-level profiling, measuring key transformer operations such as multi-head attention and general matrix multiplication.
+Since these operations are fundamental to transformer models, this approach allows APEX to support various LLMs without requiring per-model profiling.
+The profiling is performed across various context lengths, attention heads, hidden dimensions, etc.
+Additionally, the profiler measures collective communication overheads, including AllReduce, ReduceScatter, etc., across different data transfer sizes, device counts, and number of nodes.
+By covering a wide range of computation and communication configurations, the profiling results facilitate accurate estimation of execution time and energy consumption across different LLMs, device clusters, and parallel execution plans.
+Notably, profiling is a one-time cost when porting to a new cluster and can be amortized across multiple simulations.
 4. Experimental Results
@@
 Chat 73.32$\pm$148.65 189.47$\pm$174.18 1024
-4.1. Experimental Setup
-To evaluate APEX, we choose a broad range of LLMs and datasets with distinct workloads. We also experiment with different data types. We discuss the setup details below.
-Models: For evaluation, we choose four state-of-the-art LLMs: Qwen2.5-32B-Instruct (Hui et al., 2024), Llama-3.1-70B-Instruct (Dubey et al., 2024), Mistral-Large-Instruct (123B) (Jiang et al., 2023), and Mixtral 8x22B (Jiang et al., 2024); This covers a myriad of LLMs in different sizes, ranging from 32B to 123B, and also a Mixture-of-Expert (MoE) model.
-While the models use half-precision (FP16) by default, we also test cases where the model is quantized to FP8 format.
-Datasets: We prepared three sets of request traces derived from distinct datasets: a paper abstract summarization dataset (Cohan et al., 2018), a news abstract summarization dataset (Fabbri et al., 2019), and a conversational dataset, LMSYS-Chat-1M (Zheng et al., 2024).
-These datasets represent three distinct workloads.
-The paper abstract summarization (Cohan et al., 2018) exemplifies a prefill-intensive workload, which has a long context length and short generation length.
-While some LLM workloads, such as paper summarization, are prefill-intensive, some (e.g., code generation and storytelling) are generation-intensive, involving short context lengths and long generation lengths.
-To create a generation-intensive workload, we adapt the news abstract summarization dataset (Fabbri et al., 2019), which consists of short summarizations.
-We modified the dataset by appending the following prompt to each news summarization: “Please generate a long story using the provided abstract,” and use them as input.
-This prompts the LLM to produce a long story from a short abstract, resulting in a generation-intensive workload.
-The LMSYS-Chat-1M (Zheng et al., 2024) dataset collects a series of real-world conversations between users and LLMs.
-Most requests in this dataset feature short context and generation lengths, representing a lightweight conversational workload.
-For evaluation, we randomly subsampled between 512 and 1K requests from each dataset.
-To simplify terminology, we refer to these datasets as Summarization (paper abstracts), Creation (news generation), and Chat (LMSYS-Chat-1M) in the following sections.
-We assume a Poisson distribution for request arrival times.
-Details of the request traces are listed in Table 1.
-We report context lengths and generation lengths in terms of tokens, which vary for different tokenizers.
-We use the Llama-3.1 tokenizer as an example in Table 1; other tokenizers produce similar token counts.
-Hardware Platform and Serving Systems: We run APEX simulation using an Intel Xeon 6530 CPU and validated the simulation results against the actual performance of LLM serving on a GPU cluster.
-We deployed vLLM (Kwon et al., 2023) v0.6.0 to configure an LLM online server on a GPU cluster equipped with 8 Nvidia H100 SXM GPUs, each with 80 GB of GPU memory.
-Requests were sent to the server according to their arrival times.
-Since vLLM does not natively support data parallelism, we implemented data parallelism by setting up multiple vLLM servers and dispatching requests to them in a round-robin fashion.
 Table 2.
@@
 Traces Model Arrival Rate Baseline Feasible Optimal APEX Optimal
-Summarization Qwen-2.5-32B 0.25 2153.27 (1$\times$) 1137.70 (1.89$\times$) 968.26 (2.22$\times$)
-0.5 1823.85 (1$\times$) 1116.84 (1.63$\times$) 967.23 (1.89$\times$)
-Llama-3.1-70B 0.25 1998.82 (1$\times$) 1340.99 (1.49$\times$) 1175.09 (1.70$\times$)
+Summarization Llama-3.1-70B 0.25 1998.82 (1$\times$) 1340.99 (1.49$\times$) 1175.09 (1.70$\times$)
 0.5 1565.11 (1$\times$) 1225.38 (1.28$\times$) 1140.19 (1.37$\times$)
-Mistral-Large 0.25 1301.27 (1$\times$) 498.32 (2.61$\times$) 376.13 (3.46$\times$)
-0.5 1140.36 (1$\times$) 498.32 (2.29$\times$) 376.13 (3.03$\times$)
-Mixtral-8x22B 0.25 1841.90 (1$\times$) 1290.97 (1.43$\times$) 551.94 (3.34$\times$)
-0.5 1429.46 (1$\times$) 1137.87 (1.26$\times$) 552.28 (2.59$\times$)
-Creation Qwen-2.5-32B 0.25 5718.22 (1$\times$) 2432.1 (2.35$\times$) 1507.68 (3.79$\times$)
-0.5 4950.96 (1$\times$) 2421.82 (2.04$\times$) 1512.95 (3.27$\times$)
-Llama-3.1-70B 0.25 6549.85 (1$\times$) 3742.77 (1.75$\times$) 3159.96 (2.07$\times$)
-0.5 5240.78 (1$\times$) 4578.49 (1.14$\times$) 3183.67 (1.65$\times$)
-Mistral-Large 0.25 6379.09 (1$\times$) 4063.11 (1.57$\times$) 2684.37 (2.38$\times$)
-0.5 5221.52 (1$\times$) 4121.53 (1.27$\times$) 2682.05 (1.95$\times$)
-Mixtral-8x22B 0.25 2645.24 (1$\times$) 2423.82 (1.09$\times$) 1039.87 (2.54$\times$)
-0.5 1442.15 (1$\times$) 1417.75 (1.02$\times$) 1035.62 (1.39$\times$)
-Chat Qwen-2.5-32B 0.25 1251.75 (1$\times$) 798.75 (1.57$\times$) 513.38 (2.44$\times$)
-0.5 1115.01 (1$\times$) 786.08 (1.42$\times$) 510.72 (2.18$\times$)
-Llama-3.1-70B 0.25 1622.48 (1$\times$) 1118.85 (1.45$\times$) 824.43 (1.97$\times$)
+Mistral-Large 0.25 2210.05 (1$\times$) 1624.17 (1.36$\times$) 1514.19 (1.46$\times$)
+0.5 1683.00 (1$\times$) 1483.38 (1.14$\times$) 1420.98 (1.18$\times$)
+Mixtral-8x22B 0.25 2005.06 (1$\times$) 1753.79 (1.14$\times$) 859.83 (2.33$\times$)
+0.5 1618.54 (1$\times$) 1473.79 (1.10$\times$) 856.84 (1.89$\times$)
+Creation Llama-3.1-70B 0.25 4027.70 (1$\times$) 2301.54 (1.75$\times$) 1945.75 (2.07$\times$)
+0.5 1869.45 (1$\times$) 1639.87 (1.14$\times$) 1133.05 (1.65$\times$)
+Mistral-Large 0.25 3561.55 (1$\times$) 2268.50 (1.57$\times$) 1496.45 (2.38$\times$)
+0.5 2560.70 (1$\times$) 2016.30 (1.27$\times$) 1313.18 (1.95$\times$)
+Mixtral-8x22B 0.25 2645.79 (1$\times$) 2593.91 (1.02$\times$) 1496.45 (1.77$\times$)
+0.5 1393.19 (1$\times$) 1386.16 (1.01$\times$) 1035.62 (1.35$\times$)
+Chat Llama-3.1-70B 0.25 1622.48 (1$\times$) 1118.85 (1.45$\times$) 824.43 (1.97$\times$)
 0.5 1344.42 (1$\times$) 1021.09 (1.32$\times$) 808.96 (1.66$\times$)
-Mistral-Large 0.25 973.16 (1$\times$) 396.32 (2.46$\times$) 254.26 (3.83$\times$)
-0.5 890.71 (1$\times$) 391.96 (2.27$\times$) 254.26 (3.50$\times$)
-Mixtral-8x22B 0.25 1757.68 (1$\times$) 1504.88 (1.17$\times$) 397.93 (4.42$\times$)
-0.5 1406.75 (1$\times$) 1253.22 (1.12$\times$) 396.80 (3.55$\times$)
-4.2. Evaluation of APEX’s suggestion
-In this experiment, we address the following research question: Can APEX improve LLM serving performance by identifying an optimal parallel execution plan?
-For evaluation, we design multiple tasks, with each task comprising an LLM, request traces, and an arrival rate.
-We run APEX simulation for each task and compare the performance of the following three execution plans:
-  • Baseline plan: We follow the commonly-used heuristics (Miao et al., 2024; vLLM, v060), which applies tensor parallelism within the same node and pipeline parallelism across nodes.
-Since we only run on a one-node cluster, we apply tensor parallelism as the baseline plan.
+Mistral-Large 0.25 2041.91 (1$\times$) 1423.44 (1.43$\times$) 1182.75 (1.73$\times$)
+0.5 1594.74 (1$\times$) 1356.52 (1.18$\times$) 1102.20 (1.45$\times$)
+Mixtral-8x22B 0.25 1938.26 (1$\times$) 1658.22 (1.17$\times$) 575.12 (3.37$\times$)
+0.5 1507.11 (1$\times$) 1380.34 (1.09$\times$) 567.53 (2.66$\times$)
+4.1. Experimental Setup
+To evaluate APEX, we choose a broad range of LLMs and datasets with distinct workloads. We discuss the setup details below.
+Models: For evaluation, we choose four state-of-the-art LLMs: Llama-3.1-70B-Instruct, Llama-3.1-405B-Instruct (Dubey et al., 2024), Mistral-Large-Instruct (123B) (Jiang et al., 2023), and Mixtral 8x22B-Instruct (Jiang et al., 2024); This covers a myriad of LLMs in different sizes and also a Mixture-of-Expert (MoE) model.
+While the models use half-precision (FP16) by default, we also test cases where the model is quantized to FP8 format.
+Datasets: We prepared three sets of request traces derived from distinct datasets: a paper summarization dataset (Cohan et al., 2018), a news abstract summarization dataset (Fabbri et al., 2019), and a conversational dataset, LMSYS-Chat-1M (Zheng et al., 2024a).
+These datasets represent three distinct workloads.
+The paper summarization (Cohan et al., 2018) exemplifies a prefill-intensive workload, which has a long context length and short generation length.
+To create a generation-intensive workload with short context and long generation lengths, we modify the news abstract summarization dataset (Fabbri et al., 2019) by appending the following prompt: “Generate a long story using the provided abstract.” This prompts the LLM to produce a long story from a short abstract, resulting in a generation-intensive workload.
+The LMSYS-Chat-1M (Zheng et al., 2024a) dataset collects real-world conversations between users and LLMs.
+Most requests in this dataset feature short context and generation lengths, representing a lightweight conversational workload.
+We randomly subsampled between 512 and 1188 requests from each dataset to produce our request traces.
+We refer to these traces as Summarization (paper summarization), Creation (news generation), and Chat (LMSYS-Chat-1M) in the following sections.
+We assume a Poisson distribution (Consul and Jain, 1973) for request arrival times.
+Details of the request traces are listed in Table 1.
+Hardware Platform and Serving Systems: We run APEX simulation using an Intel Xeon 6530 CPU and validated the simulation results against the actual performance of LLM serving on a GPU cluster.
+We deployed vLLM (Kwon et al., 2023) v0.6.0 and SGLang v0.4.5 (Zheng et al., 2024b) to configure LLM online servers on three GPU clusters: (1) a Single-node H100 cluster equipped with 8 Nvidia H100 SXM GPUs, each with 80 GB of GPU memory, (2) a Multi-node H100 Cluster consisting of two nodes, each with 8 Nvidia H100 SXM GPUs, and (3) a H200 Cluster (single-node) that has 8 Nvidia H200 SXM GPUs.
+4.2. Evaluation of APEX’s prediction
+In this experiment, we address the following research question: Can APEX improve LLM serving performance by identifying optimal parallel execution plans?
+For evaluation, we design multiple tasks, with each task comprising an LLM, request trace, and arrival rate.
+We run APEX simulation for each task and compare the end-to-end serving latency of the following three plans:
+1. Baseline plan: We follow the commonly-used heuristics (Miao et al., 2024; vLLM, v060), which applies tensor parallelism within the same node and pipeline parallelism across nodes.
+2.
 APEX Optimal plan: APEX identifies an optimal plan for each task.
@@
 As a result, an execution plan may not be fully supported by existing LLM serving systems.
+Still, we include the results to showcase the potential performance gains that can be achieved by supporting a broader range of parallelism techniques.
+3.
 Feasible Optimal plan: This is the optimal plan identified by APEX under the constraint that only parallelism techniques supported by current LLM serving systems are used.
 Due to the reduced search space, the feasible optimal plan may achieve lower performance compared to the unconstrained APEX Optimal plan.
-While the APEX Optimal plan may not be realizable in current LLM serving systems, we include the results to showcase the potential performance gains that can be achieved by supporting a broader range of parallelism techniques.
-We show the experimental results in Table 2.
-We experiment with the three request traces and four LLMs mentioned in Section 4.1.
-We also conduct experiments using two different arrival rates for the requests, assuming a Poisson distribution.
+4.2.1. Single-Node H100 cluster
+We first evaluate APEX on a single-node cluster with 8 H100 GPUs and show the experimental results in Table 2.
+We experiment with the three request traces and three LLMs.
+We also conduct experiments using two different arrival rates for the requests, assuming a Poisson distribution (Consul and Jain, 1973).
 To explore different quantization formats, we quantize the Mistral-Large model using FP8 for the KV cache, as well as the weights and activations (i.e., W8A8).
-APEX identifies optimal parallel execution plans for both dense and sparse LLMs (i.e, MoE models), and the Feasible Optimal plans consistently deliver performance improvement over the baseline, achieving up to 2.61$\times$ speedup in terms of end-to-end serving latency.
-Furthermore, assuming cell-level data parallelism is available, the APEX Optimal Plans can achieve a speedup of up to 4.42$\times$.
-Cell-level data parallelism is particularly effective when the execution time varies among cells (e.g., MHA and MLP cells).
-This approach allows a specific cell to be parallelized further using data parallelism without requiring replication of the entire model.
-In this experiment, we demonstrate APEX is able to identify an optimal parallel execution plan that improves LLM serving performance under various LLMs and request traces.
-Figure 5. We compare the speedup predicted by APEX against the actual speedup achieved using vLLM. APEX accurately predicts the speedup of adopting an optimal execution plan over the baseline plan.
-Figure 6. APEX accurately predicts the scalability trend across varying numbers of GPUs, as measured by TPOT (ms)
-APEX identifies an optimal parallel execution plan for each task.
-The optimal plans consist of various combinations of data, pipeline, and tensor parallelism, effectively balancing the trade-offs of various factors (e.g., compute, memory, network traffic) to outperform the baseline plan, which only relies on tensor parallelism.
+APEX identifies optimal parallel execution plans for both dense and sparse LLMs (i.e, MoE models), and the Feasible Optimal plans consistently deliver performance improvement over the baseline, achieving up to 1.75$\times$ speedup in terms of end-to-end serving latency.
+Furthermore, assuming cell-level data parallelism is supported, APEX Optimal Plans achieve speedups of up to 3.37$\times$.
+In this experiment, we demonstrate APEX is able to identify an optimal parallel execution plan that improves LLM serving latency under various LLMs and request traces.
+The identified optimal plans consist of various combinations of data, pipeline, and tensor parallelism, effectively balancing the trade-offs of various factors (e.g., compute, memory, network traffic) to outperform the baseline plan, which only relies on tensor parallelism.
 While the optimal plans are different from task to task, we observe that incorporating data parallelism (DP) often yields performance benefits.
-Many of the identified optimal plans set the degree of DP to 2 or even 4.
+Many of the identified optimal plans set the degree of DP to 2.
 Existing LLM serving systems often overlook DP, assuming it is prohibitively memory-intensive.
 Instead, our results demonstrate that trading memory efficiency for reduced communication overhead can lead to performance improvements.
-We also observe that the Feasible Optimal plan of Mixtral-8x22B shows marginal improvement over the baseline on the Creation dataset.
+We also observe that the Feasible Optimal plan of Mixtral-8x22B shows only marginal improvement over the baseline on the Creation trace.
 This is due to the dominating role of memory utilization in this setup, leaving fewer trade-offs available for performance improvement.
-Specifically, Mixtral-8x22B, with its 141B parameters, heavily utilizes memory resources, while the Creation dataset’s generation-intensive nature further increases memory demands due to extensive KV caching.
+Specifically, Mixtral-8x22B, with its 141B parameters, heavily utilizes memory resources, and the Creation traces’s generation-intensive nature further increases memory demands due to extensive KV caching.
 Under these conditions, the baseline plan, which adopts pure tensor parallelism, performs well owing to its high memory efficiency.
-Nevertheless, the APEX Optimal plan still manages to deliver a 2.54$\times$ speedup by enabling cell-level data parallelism (DP), which is more memory-efficient than replicating the entire model.
+Nevertheless, the APEX Optimal plan still manages to deliver a 1.77$\times$ speedup by enabling cell-level DP.
+Table 3. End-to-End latency (sec.). APEX identifies optimal execution plans for different clusters
+Model & Cluster Traces Baseline APEX Optimal
+Llama-3.1-405B (Multi-Node H100) Summ. 4077.88 2519.12 (1.62$\times$)
+Creation 2314.25 1229.14 (1.88$\times$)
+Chat 2645.89 1710.10 (1.55$\times$)
+Llama-3.1-70B (H200) Summ. 1433.21 897.95 (1.60$\times$)
+Creation 1128.66 667.85 (1.69$\times$)
+Chat 1189.00 536.74 (2.22$\times$)
+Mistral-Large (H200) Summ. 1595.24 1174.79 (1.36$\times$)
+Creation 1389.19 776.08 (1.79$\times$)
+Chat 1515.41 961.59 (1.56$\times$)
+Mixtral-8x22B (H200) Summ. 1540.42 706.06 (2.18$\times$)
+Creation 1380.75 993.34 (1.39$\times$)
+Chat 1380.76 500.32 (2.76$\times$)
+4.2.2. Multi-Node H100 Cluster
+Next, we evaluate APEX on a two-node cluster, each node contains 8 NVIDIA H100 SXM GPUs, identical to the setup in Section 4.2.1.
+We use Llama-3.1-405B, one of the largest publicly available LLMs for evaluation on this cluster.
+Results are presented in the first row of Table 3.
+Due to space constraints, we report only the Baseline Plan and APEX Optimal Plan under an arrival rate of 0.5.
+Experimental results show that APEX consistently identifies execution plans that outperform the baseline, achieving speedups ranging from 1.55$\times$ to 1.88$\times$.
+These findings demonstrate APEX’s ability to scale effectively to multi-node clusters and extremely large models, continuing to identify performant execution plans under increased system complexity.
+4.2.3. H200 Cluster
+We also evaluate APEX on a cluster with 8 NVIDIA H200 GPUs.
+Compared to the H100 (80 GB HBM), the H200 offers the same compute and peak performance but with increased memory capacity (141 GB HBM).
+This expanded memory enables the H200 cluster to explore a larger design space, allowing for more flexible trade-offs between parallelism strategies.
+As shown in last three rows of Table 3, APEX finds simulation plans that lead to up to 2.76$\times$ speedup compared to the baseline plan.
+Notably, APEX recommends different optimal plans for the H200 and H100 clusters under the same experimental setup.
+For instance, when serving the Summarization trace with the Mistral-Large model, APEX suggests 2-way data parallelism for H100, and 4-way data parallelism plan for H200.
+While the H200 plan incurs higher memory overhead, it reduces communication overhead, making the trade-off favorable due to the H200’s larger memory capacity.
+4.2.4. Optimize Towards Energy Efficiency
+For service providers, latency is not the only metric that matters.
+Rather than minimizing inference time at all costs, providers are typically required to meet specific service-level objectives (SLOs), allowing room to optimize other metrics as long as the SLOs are satisfied.
+One critical metric is energy consumption, which directly impacts both operational costs and carbon emissions.
+APEX supports optimizing for energy consumption by profiling not only execution time but also the energy usage of operations.
+We demonstrate this capability in Table 4, which presents results for serving a Llama-3.1-70B model on a cluster with 8 H100 GPUs.
+The first rows report latency-optimal plans (i.e., the APEX Optimal Plans in Table 2), which generally yield better energy efficiency than the baseline plans due to shorter runtime.
+However, latency-optimal plans are not necessarily energy-optimal.
+We observe that energy efficiency varies across operations of different input sizes, and some operations consume disproportionately more energy; investigating the underlying cause of this phenomenon is beyond the scope of this work, and we instead focus on leveraging the collected profiling results to identify energy-efficient execution plans.
+APEX evaluates the energy consumption across various parallel execution plans and identifies the energy-optimal plan.
+Reported in the second rows of Table 4, energy-optimal plans reduce energy consumption by up to 19% compared to latency-optimal plans.
+For systems with more relaxed SLOs, service providers may further reduce energy consumption by lowering GPU frequency.
+In the third rows of Table 4, APEX identifies optimal plans under 0.8 GHz of GPU frequency, reducing energy consumption by 45% and 41% on the Summarization and Creation traces, respectively.
+While this results in increased TTFT (up to 1.9$\times$) and TPOT (up to 1.8$\times$), such increases may still be acceptable in practice, for instance, when the TPOT requirement is 50 ms.
+These results show that APEX enables service providers to make informed trade-offs between latency and energy efficiency.
+Table 4. APEX identifies plans optimized toward energy
+Traces
+Exe. Plans
+(GPU Freq., GHz)
+Energy
+TTFT
+TPOT
+Summ. Latency opt. (2.0) 4.095 788.67 32.06
+Energy opt. (2.0) 3.293 992.69 26.89
+Energy opt. (0.8) 2.265 1485.94 33.62
+Creation Latency opt. (2.0) 8.388 44.67 18.83
+Energy opt. (2.0) 8.277 54.59 20.63
+Energy opt. (0.8) 5.016 75.70 34.59
 4.3. Evaluation of Simulation Fidelity
 In this experiment, we address the following research question: Can APEX provide high-fidelity simulation for LLM serving?
-While we demonstrate in Table 2 that APEX can identify optimal parallel execution plans that enhance LLM serving performance, it is also essential to verify the accuracy of its simulation results.
+While we demonstrate in Section 4.2 that APEX can identify optimal parallel execution plans that enhance LLM serving performance, it is also essential to verify the fidelity of its simulation results.
 We evaluate the simulation fidelity from two perspectives.
@@
 We show the experimental results in Figure 6 and Figure 7.
-We set the arrival rate to 0.5 for both experiments as an example; setting the arrival rate of 0.25 yields similar results.
+Figure 6. Speedup predicted by APEX vs. actual speedup achieved. APEX accurately predicts the speedup of adopting an optimal execution plan over the baseline plan.
 For the first experiment set, we compare the predicted speedup of Feasible Optimal plans with the actual speedup achieved on the device cluster (Figure 6).
-We do not evaluate Mixtral-8x22B as vLLM currently does not support expert parallelism.
-APEX achieves accurate predictions, with an average relative error of 9.5%.
-The largest discrepancies occur with the Qwen-2.5-32B model, the smallest model in our experiment, where the relative error reaches 23%.
-This is because, for smaller models, the overhead of operations such as RMSNorm becomes relatively significant, and APEX does not account for these operations in its simulations.
-In contrast, for larger models like Llama-3.1-70B and Mistral-Large (123B), the predictions are more accurate, as the execution time is dominated by attention and feedforward cells, which are well-modeled by APEX.
-The fidelity for smaller LLMs could be improved by incorporating profiling results for additional operations.
-However, as APEX primarily focuses on optimizing performance for serving large models, which are more computationally intensive, we leave this improvement for future work.
+We use vLLM (Kwon et al., 2023) to serve Llama-3.1-70B and Mistral-Large, and use SGLang (Zheng et al., 2024b) for Mixtral-8x22B as vLLM does not support expert parallelism (EP) at the time of our experiments.
+APEX achieves accurate predictions, with an average relative error of 10.7%.
+The main discrepancy lies in the predictions for Mixtral-8x22B.
+APEX estimates that applying EP would outperform tensor parallelism (TP) due to the lower overhead of All-to-All communication in EP, compared to the AllReduce operations required in TP.
+However, when serving with SGLang, we observe that TP consistently outperforms EP across experiments.
+A potential reason is that the current implementation of EP may not be as well-optimized as TP, leading to higher-than-expected overheads during execution.
 For the second experiment set, we evaluate the time per output token (TPOT) across different numbers of GPUs using tensor parallelism.
 We show the results in Figure 7.
-The APEX prediction results are shown as blue solid lines, while the actual results are represented by red dotted lines.
+The APEX prediction results are shown as green solid lines, while the actual results are represented by red dotted lines.
 The high similarity between the two lines across all cases demonstrates that APEX accurately captures the scalability trend when scaling from 2 GPUs to 8 GPUs across various models and request traces.
@@
 The actual measured TPOT is consistently higher than the predicted value, as evident from the larger values on the right-hand Y-axis compared to the left-hand Y-axis.
-This is because APEX primarily focuses on the overhead of key operations, such as attention and feedforward, while omitting the overhead of other operations.
-Nevertheless, APEX accurately captures the relative performance differences between various parallel execution plans and degrees of parallelism, as demonstrated in Figure 5 and 6.
+This is because APEX focuses on the overhead of key operations, such as attention and feedforward, while omitting the overhead of other operations.
+Nevertheless, APEX accurately captures the relative performance differences between various parallel execution plans and degrees of parallelism, as demonstrated in Figure 6 and 7.
 This capability enables APEX to effectively evaluate and compare different execution plans, thereby identifying an optimal parallel execution plan for LLM serving.
+Figure 7. APEX accurately predicts the scalability trend across varying numbers of GPUs, as measured by TPOT (ms)
 4.4. Evaluation of Efficiency
-We evaluate the efficiency of APEX by comparing the actual execution time on the device cluster with the APEX simulation time.
+We evaluate the efficiency of APEX by comparing the APEX simulation time with the actual execution time on the device cluster.
 Evaluating all parallel execution plans across the setups in Table 2 (i.e., different models, traces, arrival rates) takes approximately 160 hours on 8 H100 GPUs.
-In contrast, the same evaluation is completed in less than 2.5 hours using APEX on CPU, making it 71$\times$ faster than the actual implementation.
-From a cost perspective, running the actual implementation would cost approximately \$currency-dollar\\$\$8,889 (based on the Azure NC40ads H100 cluster pricing), whereas the APEX simulation costs only \$currency-dollar\\$\$7.20 (assuming an Azure D64s v6 node).
-This translates to a 1234.5$\times$ cost reduction, demonstrating the significant efficiency and cost-effectiveness of APEX.
+In contrast, the same evaluation is completed in less than 2.5 hours using APEX on a CPU, making it 71$\times$ faster than the actual deployment.
+From a cost perspective, running the actual implementation would cost approximately \$currency-dollar\\$\$8,889 (based on the Azure NC40ads H100 cluster pricing), whereas the APEX simulation costs only \$currency-dollar\\$\$7.20 (assuming an Azure D64s v6 CPU node).
+This translates to a 1234.5$\times$ cost reduction, demonstrating the significant time efficiency and cost-effectiveness of APEX.
 Obtaining the operation-level profiling results to set up APEX takes approximately 40 GPU hours.
-Yet, this is a one-time cost that can be amortized for the same type of hardware device.
+Yet, this is a one-time cost that can be amortized for the same type of device cluster.
 Figure 8. APEX maintains similar simulation overhead when scaling from billion-scale to trillion-scale models
 We also evaluate the simulation overhead when scaling to larger LLMs on larger device clusters.
-We use the following example models: Qwen2.5-32B, Llama-3.1-70B, Mistral-Large (123B), and Llama-3.1-405B.
-We synthesize a trillion-scale model by scaling the Llama-3.1-70B model 16 times; this can be done by modifying values in the LLM configuration file.
-As shown in Figure 7, APEX demonstrates high scalability, maintaining a similar simulation overhead when scaling from a 32B model to a trillion-scale model.
-This efficiency is achieved by leveraging the canonical representation in Transformer IR (Section 3.3) and the repetitive structure of the transformer architecture (Section 3.7) to greatly reduce the design space.
-This scalability highlights APEX’s potential for future applications, as model sizes continue to grow and the cost of actual deployment may become even more expensive.
+We use the following example models: Qwen2.5-32B, Llama-3.1-70B, Mistral-Large, and Llama-3.1-405B.
+We synthesize a trillion-scale model by scaling the Llama-3.1-70B model 16 times; this is done by modifying values in the LLM configuration file (Huggingface, 2025).
+As shown in Figure 8, APEX demonstrates high scalability, maintaining a similar simulation overhead when scaling from a 32B to a trillion-scale model.
+This efficiency is achieved by leveraging the canonical representation of Transformer IR (Section 3.2.1) and the repetitive structure of the transformer architecture to reduce the design space.
+This scalability highlights APEX’s potential for future applications, as model sizes continue to grow and the cost of actual deployment become more expensive.
 4.5. Evaluation of Extensibility
-LLMs and their serving systems are evolving rapidly.
 While APEX supports a wide range of models, devices, etc., the simulator may still become outdated as new advancements emerge.
@@
 We evaluate APEX’s extensibility by measuring the overheads to implement a new feature.
-We use two metrics: (1) programming overhead, measured in lines of code required for a feature extension, and (2) implementation time overhead, measured in hours, including time to write and execute the necessary code or scripts.
-We reported the overhead of various types of extensions in Table 3.
-Below, we provide details of the extensions implemented for the evaluation.
-Extending to New Models APEX can effortlessly support a new LLM without additional programming or implementation time by requiring only a configuration file of the model.
-However, for LLMs containing unknown transformer cells (Section 3.3), such as a novel feedforward network, additional effort is needed.
-The primary work involves implementing the Parallel Template (Section 3.4) for the new cells.
-For evaluation, we measured the effort to support SwiGLU cells (Shazeer, 2020) used in the Llama and Qwen models.
+We use two metrics: (1) programming overhead, measured in lines of code required for a feature extension, and (2) implementation time overhead, measured in hours, including time to write and execute the code or scripts.
+We reported the overhead of various types of extensions in Table 5 and discuss the details below.
+Extending to New Models APEX can effortlessly support a new LLM without additional programming or implementation time by requiring only a configuration file (Huggingface, 2025) of the model.
+However, for LLMs containing unknown transformer cells (Section 3.2.1), such as a novel feedforward network, additional effort is needed.
+The primary work involves implementing the Parallel Template (Section 3.2.2) for the new cells.
+For evaluation, we measured the effort to support SwiGLU cells (Shazeer, 2020) used in the Llama-3.1 models.
 Extending to New Device Cluster APEX can easily adapt to new device clusters by providing the device name, memory capacity, and interconnection network (e.g., NVLink, PCIe).
-We evaluated the effort required to support a new type of GPU and an in-house AI accelerator.
-The implementation time overhead mainly involves running profiling scripts, which take 4 to 8 hours depending on the hardware’s performance.
-While extending to new devices requires a relatively long script execution time, this effort is only necessary when a new device is released.
-Extending to New Batching Mechanism APEX can also support new batching mechanisms.
-While this is not as straightforward as supporting a new model or a new device cluster, it can be done by modifying the Iteration-Level Batching Module 3.6.
-APEX adopts a vLLM-style (Kwon et al., 2023) batching by default.
-As an example, we evaluate the overhead of extending to support a Sarathi-Serve-style (Agrawal et al., 2024b) batching, which performs chunk prefilling to better interleave prefill and decode requests.
-This extension is done by adding a new variable, chunk size, to the batching module and a counter to each request to ensure all prefill chunks are completed before moving to the generation stage.
+We evaluated the effort required to support a new GPU model.
+The implementation time overhead mainly involves running profiling scripts, which take 6 to 8 hours.
+While extending to new devices requires a relatively long profiling time, this is a one-time effort that can be amortized.
+Extending to New Batching Mechanism APEX also supports new batching mechanisms.
+While this is not as straightforward as supporting a new model or a new device cluster, it can be done by modifying the Batching Module (Section 3.3).
+APEX adopts contiguous batching (Kwon et al., 2023; Zheng et al., 2024b) by default.
+We evaluate the overhead of extending to support a Sarathi-Serve-style (Agrawal et al., 2024b) batching, which performs chunked prefill to better interleave prefill and decode requests.
+This extension is done by adding a new variable, chunk size, to the Batching Module and a counter to each request to ensure all prefill chunks are completed before moving to the decode stage.
 Table 5. Evaluating the overhead of extending APEX
@@
 LLM (w/ unknown cells) 50 - 150 1-2
-Device cluster $\sim$20 4-8
+Device cluster $\sim$20 6-8
 Batching Mechansim $\sim$100 1-2
 Parallelism 50 - 200 1-2
-Extending to New Parallelism The overhead of extending APEX to support new parallelism types depends on the specific parallelism being added.
-For example, supporting a new type of intra-layer parallelism is analogous to supporting a new transformer cell, as it primarily requires adding a new Parallel Template.
-Similarly, Fully Sharded Data Parallelism (FSDP) can be integrated into APEX by extending the existing Data Parallelism implementation, incorporating additional collective communication steps, and adjusting memory usage to account for sharded model storage.
-In summary, APEX is designed to support a wide range of extensions with minimal programming and implementation overhead, enabling it to stay up-to-date with advancements in LLMs, hardware, and serving systems as they continue to evolve.
+Extending to New Parallelism The overhead of extending to new parallelisms depends on the specific parallelism being added.
+For example, supporting a new intra-layer parallelism is analogous to supporting a new transformer cell, as it requires adding a new Parallel Template.
+Similarly, Fully Sharded Data Parallelism (FSDP) can be integrated into APEX by extending the existing DP implementation, incorporating additional collective communications and adjusting memory usage for the sharded model storage.
 4.6. Beyond Identifying Optimal Execution Plan
-While APEX is primarily developed to identify optimal parallel execution plans for LLM serving systems, its high simulation fidelity makes it applicable to other use cases as well.
-Below, we present two examples of applications.
-4.6.1. APEX provides insights for building LLM serving platform
-When a new hardware device is released, service providers can use APEX to estimate the expected performance improvements from adopting the new hardware by adjusting the hardware parameters of device cluster in APEX.
-For example, the performance boost of upgrading from Nvidia A100 to H100 GPUs can be projected by scaling the compute time and data transfer time according to their relative compute power and memory bandwidth.
-Using this approach, APEX estimates that upgrading from A100 to H100 results in a 1.79$\times$ improvement in TTFT and a 1.66$\times$ speedup in TPOT; this estimation closely aligns with actual results reported in (Patel et al., 2024), which shows a 1.85$\times$ improvement in TTFT and a 1.43$\times$ speedup in TPOT.
-APEX also allows users to freely scale hardware parameters (i.e., creating synthetic hardware) to estimate the hardware upgrade required to achieve a specific performance boost in LLM serving.
-4.6.2. APEX provides insights for meeting SLOs
-In addition to minimizing end-to-end serving latency, service providers must also meet various service-level objectives (SLOs), such as maintaining a time per output token (TPOT).
-APEX can assist in achieving these SLOs.
+The high simulation fidelity makes APEX applicable to other use cases beyond identifing optimal execution plans.
+For example, APEX can help service providers meet various service-level objectives (SLOs), such as maintaining a target time per output token (TPOT).
 A common strategy to meet latency requirements is to adjust the batch size by setting a maximum batch size constraint.
 APEX begins by simulating a subset of requests to determine an upper bound for the batch size, denoted as $m$.
-It then divides this upper bound into $n$ segments and simulates the request traces with various maximum batch size constraints, ranging from $\frac{1\times m}{n}$ to $\frac{n\times m}{n}$, where $n$ is determined heuristically.
-Figure 8 illustrates two examples using the Llama-3.1-70B and Mistral-Large models on the Creation dataset.
-For simplicity, only results for pure tensor parallelism are shown.
-However, similar evaluations are performed across all parallel execution plans, resulting in numerous design points to choose from.
-Assuming a service provider aims to decrease the TPOT to meet the SLO, they can use the results to estimate the necessary adjustment to the maximum batch size constraint.
-For instance, reducing the maximum batch size constraint from 16 to 8 results in an 18% TPOT improvement for the Llama model and a 14% improvement for the Mistral model.
-However, overly restricting the maximum batch size can negatively impact end-to-end latency, as illustrated in Figure 8, where the batch size constraint is set to 4.
-Figure 8. APEX can simulate with various maximum batch size constraints and suggest adjustments to meet SLOs.
+It then divides this upper bound into $n$ segments and simulates the request traces with various batch size constraints, ranging from $\frac{1\times m}{n}$ to $\frac{n\times m}{n}$, where $n$ is determined heuristically.
+Figure 9 illustrates two examples using the Llama-3.1-70B and Mistral-Large models on the Creation trace.
+Suppose a service provider aims to reduce TPOT to meet an SLO.
+Based on APEX’s results, they can estimate the required adjustment to the batch size constraint.
+For instance, reducing the maximum batch size from 16 to 8 yields a TPOT improvement of 18% for Llama and 14% for Mistral.
+However, overly restricting the maximum batch size may negatively affect performance; APEX also captures this trade-off, as shown in Figure 9, where setting the batch size constraint to 4 leads to degraded end-to-end serving latency.
 5. Related Work
@@
 vLLM (Kwon et al., 2023) proposed PagedAttention, an efficient method to manage KV cache memory usage, enabling more concurrent request batching.
-Sarathi-Serve (Agrawal et al., 2024b) addressed prefill-decoding interference by proposing chunked prefill, a technique that breaks the prefill stage into smaller stages to enhance batching efficiency; DistServe (Zhong et al., 2024) tackled the same issue by disaggregating the prefill and decoding stages, processing each on separate device clusters optimized for their specific stages.
-Splitwise (Patel et al., 2024) also disaggregated the two stages, allocating requests to distinct device clusters to achieve higher throughput and cost-effectiveness.
-While these systems allow users to manually specify parallel execution plans (e.g., degrees of tensor and pipeline parallelism), they do not provide guidance on determining the optimal parallel execution plan.
-APEX complements these works.
-Users can utilize APEX to simulate and derive an optimal parallel execution plan tailored to their LLM serving setup, which can then be used to configure the serving system for improved performance.
-Partition-Strategy-Search Tools: Several tools have been proposed to identify an optimal configuration for LLM training.
-Calculon (Isaev et al., 2023) proposes a performance model that helps developers determine an optimal parallel execution plan for LLM training.
-vTrain (Bang et al., 2024) and ASTRA-sim (Rashidi et al., 2020) are simulators that guide users to find a system configuration that optimizes the LLM training time or the training cost.
-Nevertheless, these works cannot be applied to LLM serving due to the dynamism of iteration-level batching.
-LLMServingSim (Cho et al., 2024) and Vidur (Agrawal et al., 2024a) are the few simulation frameworks designed for LLM serving systems.
-Both perform fine-grained simulations that account for the dynamism of iteration-level batching.
-However, LLMServingSim primarily focuses on NPUs and Processing-in-Memory (PIM) architectures, as it relies on their corresponding hardware simulators to estimate execution times.
+DistServe and Splitwise (Zhong et al., 2024; Patel et al., 2024) disaggregate the prefill and decode stages, processing stage separately (e.g., using different clusters, applying different optimizations) for better serving performance.
+While these systems allow users to manually specify the degree of parallelisms, they do not provide guidance on determining these configurations.
+APEX complements these works by automatically identifying optimal plans, which can be used to configure these serving systems.
+Figure 9. APEX can simulate with various maximum batch size constraints and suggest adjustments to meet SLOs.
+Configuration Searching Tools: LLM-Pilot (Lazuka et al., 2024) proposes a predictive model to help users select the most cost-effective hardware for serving LLMs.
+DynamoLLM (Stojkovic et al., 2025) proposes a framework to automatically reconfigure the cluster to optimize the energy efficiency of LLM serving.
+However, both LLM-Pilot and DynamoLLM only consider tensor parallelism and can be further enhanced by incorporating hybrid parallelisms, as supported by APEX.
+Calculon (Isaev et al., 2023) introduces a performance model to identify optimal execution plans for LLM training, while vTrain (Bang et al., 2024) and ASTRA-sim (Rashidi et al., 2020) adopt simulator-based approaches to achieve the same goal.
+Nevertheless, these solutions are tailored to training workloads and do not address the unique challenges of LLM serving, particularly the dynamism introduced by iteration-level batching.
+LLMServingSim (Cho et al., 2024) and Vidur (Agrawal et al., 2024a) are the few frameworks designed for LLM serving.
+Yet, LLMServingSim primarily focuses on NPUs and Processing-in-Memory architectures, and it relies on their corresponding hardware simulators to estimate execution times.
 In contrast, APEX utilizes operation-level profiling data to estimate execution time and does not rely on other hardware simulators.
@@
 APEX bypasses this requirement by capturing the key operations of LLMs and utilizing pre-collected profiling results, allowing simulations to start immediately with amortized profiling costs.
-More importantly, the capabilities of both works are insufficient for simulating state-of-the-art LLM serving systems.
-Specifically, they do not support quantization techniques, their parallelism support is limited to pipeline and tensor parallelism, and they are constrained to traditional dense LLMs, lacking provisions for emerging architectures such as Mixture of Experts (MoE) models.
-While it is theoretically possible to extend these simulators to accommodate additional models and parallelism strategies, the required effort remains uncertain due to the lack of evaluations on their extensibility.
+In addition, neither LLMServingSim nor Vidur supports energy-aware optimization, a critical requirement for LLM service providers.
+They also lack support for emerging LLM serving needs, such as Mixture-of-Experts models, quantization formats, and multi-node clusters.
 6. Conlusion
 In this work, we developed APEX, an extensible and dynamism-aware simulator for LLM serving.
-We evaluated APEX across various LLMs and distinct workloads, demonstrating its high-fidelity simulation with less than 10% relative error on average.
-APEX successfully identified optimal parallel execution plans, achieving up to 4.42$\times$ speedup compared to heuristic plans.
+We evaluated APEX across various LLMs and distinct workloads, demonstrating its high-fidelity simulation with only a 10.7% relative error on the average.
+APEX successfully identified latency-optimal execution plans that achieve up to 3.37$\times$ speedup compared to heuristic plans, and also energy-optimal plans that reduce energy consumption by up to 45% compared to latency-optimal plans.
 The simulation proved highly efficient, providing a 71$\times$ speedup and 1234$\times$ greater cost-effectiveness than deployment on actual hardware.
-APEX also showcased high scalability, maintaining a similar simulation overhead when scaling from billion-scale to trillion-scale models.
-Additionally, we demonstrated APEX’s ease of extensibility to accommodate new models, device clusters, and more.
+APEX maintained a similar simulation overhead when scaling from billion-scale to trillion-scale models.
+While APEX already supports a broad range of LLMs, we further demonstrated its ease of extensibility to accommodate new models, device clusters, batching algorithms, and parallelisms.
 In the future, we plan to extend APEX to support multimodal LLMs, which will involve incorporating parallel execution plans for encoders handling modalities such as vision and audio.